### v1

# A unified framework for geometry-independent operator learning in cardiac electrophysiology simulations

Abstract
Accurate maps of atrial electrical activation are essential for personalised treatment of arrhythmias, yet biophysically detailed simulations remain computationally intensive for real-time clinical use or population-scale analyses. Here we introduce a geometry-independent operator-learning framework that predicts local activation time (LAT) fields across diverse left atrial anatomies with near-instantaneous inference. We generated a dataset of 308,700 simulations using a GPU-accelerated electrophysiology solver, systematically varying multiple pacing sites and physiologically varied conduction properties across 147 patient-specific geometries derived from two independent clinical cohorts. All anatomical and functional data are expressed in a Universal Atrium Coordinate system, providing a consistent representation that decouples electrophysiological patterns from mesh topology. Within this coordinate space, we designed a neural operator with a vision-transformer backbone to learn the mapping from structural and electrophysiological inputs to LAT fields. With a mean prediction error of 5.1 ms over a 455 ms maximum simulation time, the model outperforms established operator-learning approaches and performs inference in 0.12 ms per sample. Our framework establishes a general strategy for learning domain-invariant biophysical mappings across variable anatomical domains and enables integration of computational electrophysiology into real-time and large-scale clinical workflows.
1 Introduction
Atrial fibrillation (AF) is the most common sustained cardiac arrhythmia, affecting around 46.3 million individuals worldwide [kornej2020epidemiology], and represents a major cause of stroke [cheng2024global], heart failure [li2025global], and sudden cardiac death [lai2025causes]. Computational simulations have emerged as powerful tools in AF research and clinical trials, offering insights into arrhythmia mechanisms, enabling the prediction of ablation outcomes, and personalised therapeutic strategies [heijman2021computational, boyle2019computationally, roney2021time, krogh2012effects]. Many of these applications rely on calibrating patient-specific biophysical models to local activation time (LAT) maps, which capture the spatiotemporal propagation of electrical wavefronts across the myocardium. This calibration typically involves solving large ensembles of high-fidelity finite element method (FEM) simulations that numerically solve the governing partial differential equations (PDEs) of cardiac electrophysiology on anatomically accurate patient-specific geometries [goktepe2009computational, rocha2010macro].
Although highly accurate, FEM workflows are computationally expensive and often take hours to days to solve, making them impractical for intraoperative use. Consequently, patient-specific models are usually prepared in advance from pre-procedure MRI or CT scans that define anatomy and tissue properties. In many clinical scenarios, however, such imaging is unavailable, and the only anatomical and electrophysiological data come from intraoperative electro-anatomical mapping (EAM). While EAM-derived LATs can be used for retrospective calibration [jaffery2024review], the turnaround time of FEM-based methods prevents their use for real-time decision-making during ablation.
Recent advances in machine learning have opened new opportunities to accelerate the numerical solution of cardiac electrophysiology models. Physics-informed neural networks (PINNs) have been proposed as a physics-constrained alternative to purely data-driven surrogates, embedding the governing PDE residuals directly into the loss function [pinn]. They are typically trained for a single PDE configuration, defined by a fixed geometry, parameter set, and boundary conditions [kovachki2023neural]. However, they do not naturally generalise to new configurations without retraining: if the domain shape, material parameters, or boundary conditions change, the PINN must be re-trained from scratch or at least fine-tuned.
Neural operators directly address this limitation by learning a mapping between infinite-dimensional function spaces rather than a single instance of a PDE [kovachki2023neural]. Foundational neural operators such as the Deep Operator Network (DeepONet) [don] and the Fourier Neural Operator (FNO) [fno] have demonstrated strong performance on canonical problems defined over structured domains. By learning this operator mapping, they can efficiently predict solution fields for new inputs, even for unseen geometries or different discretisation resolutions, without the need for time-consuming retraining. Applying neural operators to complex real-world problems, such as predicting cardiac activation maps, presents significant challenges. First, these aforementioned neural operator frameworks are built to handle data on structured, canonical domains, making them ill-suited for diverse and irregular geometries [geofno, gino]. Second, training a reliable neural operator is contingent on the availability of large, high-quality datasets of diverse simulation results [ahmad2024diffeomorphic], and generating such datasets is a major challenge [trayanova2021machine].
Several approaches have been proposed to address the challenge of anatomical variability. Diffeomorphic Mapping Operator Learning (DIMON) [yin2024scalable] uses diffeomorphic transformations to map individual geometries to a shared, standardised space. This allows a neural operator to learn latent solutions on the reference domain, with the original shape encoded as an auxiliary input. While effective for relatively simple geometries, such as the semi-spherical left ventricle (LV), this strategy faces substantial difficulties when extended to anatomies with more intricate topologies, such as the multi-holed structure of the left atrium, where constructing topologically consistent reference domains is non-trivial. Similarly, methods like GeoFNO require a diffeomorphic mapping to simple domains (e.g., cubes or tori) where the Fourier transform can be applied [geofno]; however, the topologically complex 3D geometries of the human left atrium, with multiple openings, make such mappings impossible [loeffler2024graph]. Graph-based methods like the Graph Fourier Neural Kernel (G-FuNK) [loeffler2024graph] learn directly on mesh graphs, avoiding domain mapping. However, their spectral bases are sensitive to mesh quality and topology, reducing robustness and generalisability in heterogeneous clinical datasets.
Here, we introduce a method that overcomes the dependency of neural operators on structured domains for electrophysiology simulations. Our approach projects inputs and outputs onto a common, anatomically standardised reference using the Universal Atrial Coordinates (UAC) system [uac]. The UAC defines a mesh-independent coordinate space for the left atrium (LA), preserving physiologically relevant spatial relationships while unifying geometry and topology across patients. This representation enables neural operators such as DeepONet and FNO to consistently interpret spatial features, free from irregular mesh geometries or topological inconsistencies.
To support large-scale training, we developed a GPU-accelerated FEM solver, drastically reducing the time required to compute LAT maps and enabling the generation of a high-fidelity, comprehensive dataset. Using this dataset, we train a neural operator to approximate the underlying biophysical mapping, achieving near real-time inference without compromising physiological accuracy. Our model, an encoder-decoder architecture with a Vision Transformer (ViT) [vit] backbone, predicts full-field LAT maps from patient-specific anatomical features, conductivity, and pacing sites. It achieves a mean absolute error of $5.1\pm 0.2$ ms relative to FEM simulations, exhibiting high pointwise accuracy, spatial smoothness, and robust generalisation across independent clinical cohorts, overcoming the domain shift inherent in multi-centre medical data.
By integrating GPU-accelerated simulation and neural operator training into a unified framework, our approach addresses both the computational bottleneck of large-scale electrophysiology simulations and the geometric variability of patient-specific atrial anatomies. It enables accurate prediction of full-field LAT maps in just 0.12 ms per sample on a GPU, offering a pathway toward rapid, patient-specific, physiology-informed decision support for AF interventions.
2 Results
2.1 Large-scale EP simulations on GPUs
We generated a comprehensive set of biophysically detailed EP simulations using a custom GPU-accelerated FEM solver applied to two independent cohorts of AF patients (Cohort A: $n=100$; Cohort B: $n=47$). This process constructed two large-scale datasets, Dataset A and Dataset B, respectively (see Methods). To ensure robustness across diverse anatomical and physiological conditions, for each of the $147$ patient-specific LA, we simulated LAT maps from seven distinct pacing sites across the LA. At each site, we sampled $300$ combinations of the key conduction parameters, the longitudinal conductivity ($\sigma_{l}$) and the anisotropy ratio ($\sigma_{l}/\sigma_{t}$), using Latin hypercube sampling [mckay2000comparison] to cover a wide range of clinically relevant tissue properties. This yields a total of 308,700 simulated LAT maps, establishing one of the largest computational datasets of its kind for training neural operators.
Figure 1: Construction and characterisation of large-scale electrophysiology datasets generated from biophysically detailed simulations. a, Example local activation time (LAT) map from a 3D left atrium (LA) electrophysiology simulation with coronary sinus pacing, together with its projection onto the standardised Universal Atrial Coordinate (UAC) domain. b, UMAP embedding of all UAC-projected LAT maps from Datasets A and B, showing seven pacing-site clusters per cohort and systematic displacement between corresponding clusters, indicating a pronounced domain shift between the two datasets. c, Distribution of maximum LAT across pacing locations in Dataset A, demonstrating linear dependence on LA surface area and sensitivity to conductivity. d, Corresponding analysis for Dataset B, showing the same physiological trends despite cohort-level differences.
All simulated LAT maps were subsequently projected onto the standardised UAC domain to obtain a consistent, geometry-independent representation. Fig. 1a illustrates an LAT map on the 3D atrial surface together with its corresponding 2D UAC projection, with the coronary sinus pacing site producing full atrial activation within 204 ms.
To assess potential cohort-level differences, we performed dimensionality reduction on the full set of UAC-projected LAT maps. High-dimensional LAT fields were first denoised using principal component analysis (PCA), and the resulting principal components were subsequently embedded using Uniform Manifold Approximation and Projection (UMAP) [healy2024uniform]. The 2D embedding (Fig. 1b) reveals a clear separation between the two datasets. Within each cohort, seven well-defined clusters emerge, each corresponding to one of the pacing sites. Clusters associated with the same pacing location in Datasets A and B do not overlap but instead form displaced pairs in the latent space. This consistent displacement provides visual evidence of a domain shift between the cohorts, arising not from physiological differences, but from systematic variations in anatomical distributions, acquisition protocols, or preprocessing pipelines. This domain shift has a measurable impact on cross-domain generalisation performance.
We next quantified the primary determinants of total activation time across all simulations. Three factors exert the strongest influence: pacing site, LA surface area, and conduction velocities. The maximum LAT varies across pacing locations in Dataset A (Fig. 1c) and Dataset B (Fig. 1d). Across both datasets, maximum LAT increases approximately linearly with LA surface area and decreases with increasing conductivity values. These consistent trends confirm that surface area and conductivity jointly modulate global activation timing, motivating their inclusion as key anatomical features in the subsequent analyses.
2.2 Input feature sensitivity analysis
Figure 2: Comprehensive evaluation of model performance for LAT map prediction. a, Our model performance from feature sensitivity tests on Dataset A, demonstrating the contribution of each input feature. b, Evaluation of the impact of two regularisation terms, namely total variance (TV) and Laplacian, on LAT map prediction on Dataset A. c, Performance comparison between our proposed model and established deep learning architectures, including DeepONet, Fourier Neural Operator (FNO), Wavelet Neural Operator (WNO), U-Net, and ResNet on Dataset A. d, Spatial distribution of the averaged LAT prediction error (in ms) across five test cases in a representative validation fold. White circles highlight regions with the highest prediction errors, and adjacent numbers show the average error within those specific regions. e, Performance comparison between Single Pacing Location Models and a Unified Model across seven pacing sites. f, Cross-domain performance comparison on two distinct datasets. The notation $X\rightarrow Y$ indicates that the model was trained on source dataset $X$ and evaluated on target dataset $Y$.
We evaluated the contribution of each anatomical input modality by removing spatial coordinates, fibre orientation, and surface area features in turn, and retraining the model under each condition. Performance was assessed using two complementary metrics: Mean Absolute Error (MAE), which quantifies the numerical accuracy (error in milliseconds), and the Structural Similarity Index Measure (SSIM) [wang2004image], which measures the spatial fidelity and structural quality of the predicted activation maps. All analyses in this section were performed using 5-fold cross-validation. Unless stated otherwise, experiments were conducted on Dataset A, with cross-domain evaluations involving both Datasets A and B.
The results in Fig. 2a show that incorporating all three feature types achieves accurate and spatially smooth LAT predictions (MAE $6.09\pm 0.26$ ms; SSIM $0.963\pm 0.001$). Removing the surface area feature results in a pronounced degradation in performance (MAE $9.88\pm 0.81$ ms), indicating that global anatomical scaling is an essential contextual signal for correct wavefront organisation. Similarly, excluding spatial coordinates reduces both numerical accuracy and structural similarity (MAE $9.04\pm 0.44$ ms; SSIM $0.948\pm 0.002$), demonstrating that explicit localisation in the UAC domain is critical for resolving conduction pathways.
In contrast, removing fibre orientation led to improved performance across both metrics (MAE $5.15\pm 0.25$ ms; SSIM $0.978\pm 0.002$). This suggests that the fibre orientation information, as currently represented, may introduce noise or redundancy relative to the geometric cues already encoded in the coordinate and surface area channels. We discuss this observation further in the discussion section. Accordingly, fibre orientation features were excluded from all subsequent experiments.
2.3 Evaluating the effect of regularisation
To assess the impact of regularisations on LAT map reconstruction, we evaluated the contribution of two common spatial regularisation terms, total variation and second-order Laplacian loss, when added individually or in combination. These regularizers aim to encourage spatial smoothness and suppress high-frequency artefacts in the predicted LAT maps.
To evaluate the impact of different loss formulations on model performance, we conducted systematic experiments by varying the regularisation coefficient $\lambda\in\{5\times 10^{-2},3\times 10^{-2},1\times 10^{-2},5\times 10^{-3},1\times 10^{-3}\}$ in the loss function.
For each fold in 5-fold cross-validation, the model achieving the best performance on the validation set was retained, and the corresponding metrics were recorded. The performance for a given regularisation weight $\lambda$ was derived by averaging these metrics across the five folds. To probe the effect of each regularisation configuration, such as TV alone or the combination of TV and Laplacian, we further averaged the fold-averaged metrics across all tested $\lambda$ values (Fig. 2b).
Among all evaluated configurations, the $H^{1}$ loss combined with both TV and Laplacian regularisations achieved the best overall performance, with a mean MAE of $5.16$ ms and a mean SSIM of $0.978$. This demonstrates that combining both regularisation terms effectively enhances structural fidelity while preserving numerical accuracy. The addition of TV or Laplacian individually yielded smaller improvements, suggesting that their combined effect is necessary to fully capture the relevant spatial structures.
2.4 Benchmarking model performance
We benchmarked a range of neural operator frameworks and deep-learning architectures capable of modelling cardiac activation patterns from multimodal inputs. For each model, we selected three primary hyperparameters governing capacity and constructed six configurations by varying these across representative ranges, ensuring that the comparison is both systematic and approximately capacity-matched across architectures (Fig. 2c). Supplementary Section 2 provides further details on architecture design and training procedures.
We first evaluated operator learning approaches. While the Fourier Neural Operator (FNO) [fno] is theoretically well-suited for learning global mappings in function space, its empirical performance was moderate (MAE $10.14\pm 0.07$ and SSIM $0.938\pm 0.005$), indicating potential challenges in handling heterogeneous and noisy physiological data. The Wavelet Neural Operator (WNO) [wno] exhibited the weakest performance among evaluated models (MAE $10.08\pm 0.28$ and SSIM $0.938\pm 0.007$, likely reflecting sensitivity to measurement noise and anatomical variability inherent in real-world datasets. The DeepONet framework, using our proposed ViT-based architecture for both branch and trunk networks, achieved strong performance (MAE $7.12\pm 0.37$ ms; SSIM $0.963\pm 0.005$).
Among convolutional baselines, the U-Net architecture [unet] outperformed the simpler ResNet model [resnet]. The U-Net’s multi-scale encoder-decoder structure, with its strong inductive bias for spatial localisation, yielded a balanced performance of MAE $8.48\pm 0.28$ ms and SSIM $0.965\pm 0.004$. However, these convolutional models were fundamentally limited in their ability to capture global pacing effects and long-range spatial dependencies
Our model surpassed all baselines, achieving the best results with an MAE of $5.15\pm 0.25\text{ ms}$ and SSIM of $0.978\pm 0.002$. This confirms the model’s superior capability to synthesise heterogeneous inputs into spatially coherent, smooth, and highly accurate LAT map predictions.
While MAE and SSIM provide global measures of model performance, they do not reveal where errors concentrate spatially. We therefore visualised the prediction error distributions across the UAC domain. Fig. 2d displays the spatial distribution of prediction error on LAT maps for six test cases, each averaged over seven pacing locations and 300 combinations of conductivity velocities. Across all six test cases, a visually consistent and highly coherent source of localised prediction error is observed near the horizontal central line of the UAC map. This localised band of elevated error is the most prominent spatial feature in the error distribution plots. Beyond this central band, the prediction errors generally remain low and diffuse across the rest of the UAC domain, indicating high model fidelity in most tissue regions.
Beyond predictive accuracy, we further assessed the computational efficiency of our model to evaluate its suitability for real-time applications. Inference on batches of 1000 samples was performed on an NVIDIA RTX 5090 GPU in evaluation mode with CUDA synchronisation to ensure precise wall-clock timing. The model achieved an average inference time of $122.23$ ms per batch, or approximately $0.12$ ms per sample. This represents a five to six orders of magnitude acceleration compared with conventional finite-element electrophysiology solvers, which require approximately one minute per simulation with our solver in this study. Such efficiency enables near real-time prediction of LAT maps across diverse atrial anatomies and pacing configurations, highlighting the potential of our model for integration into time-critical clinical workflows and large-scale population studies.
2.5 Model generalisation across pacing sites
Understanding whether a single model can generalise across diverse pacing conditions is critical for building scalable and clinically useful predictors of atrial activation, as it allows us to determine whether it can match or exceed the performance of individually trained models. Success in this setting would indicate that the model captures shared electrophysiological structure, enabling robust predictions without prior knowledge of stimulation location.
We thus assessed the predictive performance of two modelling paradigms for LAT map reconstruction: (i) Single Pacing Location Models, trained independently for each of seven pacing sites, and (ii) a Unified Model, trained jointly across all sites to learn a shared inductive bias over anatomical and physiological structure. Across all pacing conditions, the Unified Model yields consistent and statistically significant improvements over its independently trained counterparts. Specifically, it achieves a 16.5% reduction in mean MAE from 6.3 ms to 5.2 ms (Fig. 2e, left), demonstrating enhanced predictive accuracy across heterogeneous inputs. The model also delivers a 1.67% increase in average SSIM from 0.960 to 0.976 (Fig. 2e, right), reflecting superior preservation of spatial structure and clinically relevant activation patterns.
2.6 The challenge of latent data heterogeneity
Our initial experiments confirmed the strong performance of our model on a single dataset (Dataset A), demonstrating its ability to capture anatomical features relevant to the target task within a controlled domain. However, high performance within the same dataset does not guarantee translational reliability. Here, we distinguish between internal performance, which measures the model’s accuracy when trained and tested on the same dataset, and cross-domain performance, which measures generalisation to an independent dataset. Evaluating both is critical to determine whether the model has learned biologically meaningful features or relies on dataset-specific cues.
Cross-domain generalisation was evaluated on two levels. First, we measured model performance when trained and tested on the same dataset (internal performance: $A\rightarrow A$ and $B\rightarrow B$) compared with performance when trained on one dataset and tested on the other (cross-domain performance: $A\rightarrow B$ and $B\rightarrow A$). Second, we assessed whether including data from an additional domain ($(A+B)\rightarrow A$, $(A+B)\rightarrow B$, and $(A+B)\rightarrow(A+B)$) during training improves generalisation to the original dataset, using internal performance as a baseline.
Each evaluation was performed on models trained with 18 combinations of hyperparameters to ensure statistical significance, with results highlighting consistent trends (Fig. 2f; full results in Supplementary Section 3). The strong internal performance confirms that the UAC projection effectively separates anatomical representation from patient-specific shape. In contrast, the pronounced drop in cross-domain performance provides clear evidence of latent bias. The model relies on image-specific cues that mislead predictions when applied to a new domain. This finding highlights a second-order challenge in translational modelling. Models are highly sensitive to systematic variations in data representation and annotation rather than true anatomical differences.
Integrating heterogeneous datasets during training mitigates these domain-specific biases. Training on the combined datasets $(A+B)$ encourages the model to focus on consistent anatomical features, ignoring spurious correlations tied to individual acquisition or annotation pipelines. Across both datasets, MAE decreased and SSIM increased in comparison with the internal performance, demonstrating that exposure to independent cohorts acts as a form of regularisation. For completeness, we also evaluated the model trained and tested on the combined datasets $(A+B)\rightarrow(A+B)$, obtaining $(5.77\pm 0.16)$ for MAE and $(0.9734\pm 0.003)$ for SSIM.
3 Discussion
The increasing complexity of modern biomedical simulations, particularly in fields like computational cardiology [brown2025cardiac], has created a growing need for methods that can bridge the gap between high-fidelity mechanistic models and the demands of real-time clinical applications. The results demonstrate that our proposed framework can accurately and robustly learn the complex, nonlinear mapping from a patient’s anatomy and electrophysiological parameters to their full-field LAT map, achieving high fidelity to biophysical simulations in milliseconds. This represents an advancement over conventional approaches and a crucial step toward enabling in-procedure, real-time, personalised cardiac electrophysiology analysis.
3.1 Geometry-independent neural operator learning
Our approach demonstrates that a well-designed coordinate system is a powerful tool for bridging the gap between irregular, real-world data and the structured-domain requirements of neural networks, allowing a single neural operator to learn a geometry-independent mapping. By leveraging the UAC system, we established a canonical representation that enabled neural operators to robustly capture the complex, nonlinear physics of cardiac conduction [hao2023gnot, li2023transformer, shih2025transformers, ovadia2024vito]. A key finding is the model’s low MAE of $(5.15\pm 0.25)$ ms. This level of accuracy is significant in a clinical context, as it is well within the typical measurement noise and physiological variability observed in cardiac electroanatomic mapping systems [gaeta2021high] (typically on the order of 5–10 ms [narayan2024advanced]). This demonstrates that our model’s predictions are not only numerically accurate but also fall within a clinically acceptable margin, making the approach viable for non-invasive patient stratification, procedure planning, and guiding ablation therapies.
A central tenet of our work is the understanding that the success of a neural operator relies on the availability of a large, diverse, and high-quality training dataset. By developing a GPU-accelerated FEM solver, we were able to generate a large simulation dataset across diverse patient-specific left atrial anatomies, providing a robust foundation for learning generalizable principles of cardiac electrophysiology. This data generation pipeline is a critical component of our framework, as it bypasses the prohibitive computational bottleneck that has historically limited the application of data-driven methods in this domain.
Beyond supporting our framework, this large, curated dataset provides a foundation for future data-driven studies in cardiac electrophysiology. Its scale, diversity, and consistency enable predictive models to generalise across anatomies and protocols, a capability previously limited by data scarcity. For comparison, G-FuNK [loeffler2024graph], trained on 25 left atrial meshes from atrial fibrillation patients, concluded that additional geometries could improve performance. DIMON [yin2024scalable] used 1,006 left ventricular patient data, and each was simulated at seven pacing sites, totalling 7,042 simulations. In contrast, our simulations preserve these locations while also varying tissue conductivity, allowing neural operators to capture electrophysiological responses across diverse stimulation patterns and heterogeneous, patient-specific left atrial anatomies. Overall, our dataset spans 308,700 simulations, offering a far richer foundation for training and benchmarking state-of-the-art neural operators.
3.2 Comparative model analysis
Our model achieves the best performance, which can be attributed to two key factors. First, the self-attention mechanism enables global context modelling [vaswani2017attention], allowing the network to capture long-range dependencies such as conduction wavefront spread from distant pacing sites. Second, by using convolutional patch embeddings and transposed convolution decoding, ViT maintains strong spatial localisation [wang2022convolutional], which is critical for resolving the high-frequency components of activation fronts. This balance between global awareness and local detail explains ViT’s superior pointwise accuracy and its strong preservation of spatial gradients.
We benchmarked our approach against several state-of-the-art neural operators and two different CNN architectures. Our model’s performance surpassed these baselines for reasons rooted in their fundamental design. For instance, DeepONet, which in this work employs our model for its branch or trunk networks, leverages the same global context modelling advantages as our model. This enables DeepONet to perform reasonably well, particularly in SSIM and $H^{1}$ loss, by effectively capturing smooth functional mappings and spatial dependencies. However, despite incorporating ViT components, DeepONet’s architecture focuses on decomposing input-output relationships into branch and trunk networks, which may limit its end-to-end spatial locality learning compared to the full ViT model. Consequently, its MAE performance lags behind ViT, reflecting some difficulty in the precise localisation of rapid feature changes critical for accurate cardiac wavefront propagation. In contrast, FNOs are designed to learn global mappings in function space using spectral convolutions. While FNOs can capture global trends efficiently, their reliance on low-frequency Fourier modes makes them ill-suited for capturing sharp transitions and local gradient variations common in LAT maps, particularly near pacing initiation sites. This is reflected in their relatively poor $H^{1}$ loss and MAE, despite modest SSIM performance. WNOs attempt to address this by incorporating multi-resolution representations, theoretically improving sensitivity to localised patterns. However, WNO’s performance degrades substantially in this setting, likely due to instability in wavelet decomposition on structured but irregular biomedical domains and sensitivity to small-scale alignment errors, leading to high variance in $H^{1}$ loss.
Among the convolutional baselines, U-Net delivers strong SSIM, indicating that it generates visually plausible LAT maps. Its hierarchical multi-scale architecture with skip connections allows for reconstruction of structural details. However, convolutional models like U-Net and ResNet-style CNN architectures are inherently local in receptive field, making it difficult to capture global pacing effects or subtle long-range interactions without very deep networks or large kernels [luo2016understanding].
3.3 Toward domain-invariant generalisation
Prior studies have evaluated model performance primarily under variations in pacing site [yin2024scalable]. In contrast, a central contribution of our work is demonstrating that a single model can generalise across both diverse pacing locations and simulation parameters. This capability is essential for clinical applicability, since the pacing site is often variable or unknown in practice. Our results also show that the unified model, trained jointly across all seven pacing sites, consistently and significantly outperforms models trained for individual pacing locations. The unified model achieves a 16.5% reduction in mean MAE, and a 1.6% increase in average SSIM. This demonstrates that our model effectively extracts generalizable electrophysiological principles that are independent of the specific pacing location, making it a robust and scalable solution for real-world clinical applications where prior knowledge of the stimulation site is unavailable.
Our cross-domain experiments reveal a critical insight. The apparent limitation of single-domain models is not a shortcoming of the ViT architecture, but a reflection of latent heterogeneity in clinical imaging data. By explicitly combining datasets from distinct acquisition and annotation protocols, we demonstrate that the model naturally discards protocol-specific shortcuts and focuses on the fundamental anatomical structure. This approach not only rescues performance on external cohorts but also slightly improves internal predictions.
This finding establishes a generalizable strategy for achieving domain-invariant medical image analysis. Our UAC projection provides the necessary geometry-invariant foundation, while training on heterogeneous data acts as a regularization method in feature space, compelling the model to learn only the biophysically relevant structures common to both domains and emphasizing that true translational power derives from integrating diverse, multi-centre data rather than from isolated single-site optimisation.
3.4 Limitations and future directions
Despite its strong performance, the present study has limitations. The observation that removing fibre orientation features improved model performance was initially counterintuitive. Our analysis suggests two primary reasons for this finding. First, the current representation of fibre orientation may contain inherent noise or inaccuracies that interfere with the model’s ability to learn the true underlying electrophysiological dynamics. Second, the model appears to be leveraging the spatial and geometric priors provided by the coordinates and surface area features more effectively. These geometric inputs may implicitly contain sufficient information about the overall anatomical structure, making the explicit fibre orientation data redundant or even counterproductive in its current form.
Beyond the feature representation, limitations also arise from the use of the UAC system. The UAC system is designed to map the closed, complex 3D left atrial surface onto a flat, 2D grid. To achieve this flattening, a continuous line must be cut on the 3D surface. Depending on the chosen coordinate orientation, this seam manifests as a central line on the 2D map. This seam creates a spatial discontinuity for the learning model, as points that are physically adjacent on the 3D surface are severed across this line in the 2D domain. Consequently, the neural operator struggles to accurately interpolate the LAT function across this boundary, resulting in a consistent band of large error centred precisely along this necessary central UAC seam.
The additional large localised errors, which are more distributed, are plausibly linked to areas of extreme anatomical complexity and steep electrophysiological gradients. These clusters primarily occur around the highly patient-specific structures like the PV junctions and LAA. The UAC projection cannot fully remove the residual geometric variability present in these complex regions, increasing the model’s uncertainty. Furthermore, activation wavefronts often exhibit the steepest LAT gradients in these areas. The model is challenged to precisely localise and reproduce these sharp functional variations across diverse anatomies, leading to the small, perceptible clusters of higher error in these critical physiological locations. These difficulties suggest that while our model excels at capturing global activation patterns, the fine-grained, localised details influenced by complex anatomical geometries remain a challenging frontier. Future work could explore incorporating a more fine-grained representation of these critical regions to further enhance local accuracy.
Finally, the framework was trained exclusively on high-fidelity simulated datasets. Although these simulations spanned a wide range of anatomies and electrophysiological parameters, direct validation against clinical electroanatomic mapping data is essential to confirm robustness under real-world conditions. The current implementation focuses on the left atrium; extending the approach to whole-heart or biventricular modelling will require adapting the anatomical projection and scaling the data generation process. Future directions include combining hybrid simulation–clinical datasets, refining anatomical embeddings, and exploring adaptive spatial resolutions to capture fine-scale electrophysiological detail without compromising inference speed.
4 Methods
We developed a unified, scalable computational framework that overcomes the computational bottlenecks of conventional EP simulations (Fig. 3). This framework is designed to circumvent the computational bottleneck inherent in conventional EP simulations. It leverages a diverse, high-fidelity dataset generated by a custom high-performance FEM solver to enable the training of a neural operator that can predict LAT maps with accuracy comparable to ground truth, but within milliseconds.
Figure 3: The overview of our proposed computational framework for operator learning for electrophysiology (EP) simulations. a, Preparation of left atrium (LA) mesh, mapping the 3D vertex coordinates defining the mesh to a 2D grid by Universal Atrial Coordinates (UAC), and identifying the pacing locations. b, The application of our GPU-based FEM solver to produce a large amount of simulated local activation time (LAT) maps, which are also projected to 2D grids by UAC. c, The training of neural operators using the structured simulation data.
4.1 Image data processing
4.1.1 Datasets description
The full dataset comprised 147 patient-specific left LA anatomies derived from two large-scale clinical cohorts. These anatomies served both as inputs to the EP simulations and as training and evaluation data for the neural operator models. To support both internal evaluation and cross-domain generalisation analysis, the 147 anatomies were partitioned into two distinct cohorts (Cohort A and Cohort B).
Cohort A, consisting of 100 AF LA samples (43 paroxysmal, 41 persistent, and 16 long-standing persistent), was obtained from the AF Recurrence Cohort [roney2022predicting], which was recruited in London, UK. LA geometries were segmented from late gadolinium-enhanced cardiac magnetic resonance (LGE-CMR) images acquired at end-diastole, yielding three-dimensional anatomical surfaces including the pulmonary veins and mitral annulus. Fibrotic tissue distribution was quantified using standard LGE intensity thresholding, providing patient-specific fibrosis maps that informed the EP simulations. This cohort served as the controlled domain for model development and internal validation.
Cohort B, containing 47 LA samples from AF patients recruited from Amsterdam, The Netherlands, was sourced from an independent clinical study [lee2025regional]. The LA geometries in this cohort underwent the same segmentation and fibrosis quantification pipeline as those in Cohort A to maintain structural consistency.
As LGE-CMR does not resolve myofibre orientation, fibre direction vectors for anisotropic conduction were generated for all 147 anatomies using a model based on high-resolution ex-vivo Diffusion Tensor Magnetic Resonance Imaging (DTMRI) data [pashakhanloo2016myofiber]. These fibre fields were subsequently incorporated into an atrial fibre atlas and projected onto each patient-specific geometry using the UAC system, as described by [roney2021constructing].
Notably, Cohorts A and B were derived from two distinct and independent studies conducted by separate research groups operating within independent healthcare systems. This independence in data acquisition and analysis supports robust cross-domain evaluation. Moreover, despite adhering to equivalent segmentation and processing workflows [solis2023evaluation, razeghi2020cemrgapp], the two cohorts exhibit meaningful domain shifts. Variations arise from MRI acquisition parameters (scanner vendor, pulse sequences, and field strength), from different practices in anatomical trimming (particularly of pulmonary veins and the LA appendage), and from inter-operator differences in anatomical landmarking and boundary definitions. These protocol-level and anatomical choices shape the resulting geometries and fibrosis patterns, such that models trained exclusively on Cohort A typically encounter Cohort B as an out-of-distribution domain. These differences are not physiological changes. These differences are not physiological. They are artefacts of data preparation, yet they change the geometric domain on which the operator is defined.
4.1.2 UAC projection
We projected the spatial coordinates and fibre directions on the mesh, along with the resulting LAT map, to structured grids of fixed size by UAC. The UAC system [uac] defines a bijective transformation
$\Phi:\mathcal{M}\rightarrow[0,1]^{2},$
where $\mathcal{M}\subset\mathbb{R}^{3}$ denotes the atrial surface manifold, and $0,1]^{2}$ is a normalised two-dimensional coordinate space. The map $\Phi$ is constructed using anatomical landmarks, typically the pulmonary vein junctions and the mitral annulus, as anchor points to define physiologically meaningful coordinates. These coordinates correspond to a geodesic distance-based unfolding of the atrial surface, such that anatomical regions (e.g., anterior wall, posterior wall, roof) consistently map to aligned regions of the 2D domain across patients.
The standardised representation allows to map topological entities from different anatomy to corresponding regions. Let $f:\mathcal{M}\rightarrow\mathbb{R}$ be a scalar function defined on the atrial mesh (e.g., LAT values), and let $(\alpha_{i},\beta_{i})=\Phi(\mathbf{x}_{i})$ be the UAC coordinates of a mesh vertex $\mathbf{x}_{i}$. We then construct a function $\tilde{f}:[0,1]^{2}\rightarrow\mathbb{R}$ via spatial interpolation:
$\tilde{f}(\alpha,\beta)=\mathcal{I}\left(\left\{(\alpha_{i},\beta_{i},f_{i})\right\}_{i=1}^{N}\right),$
where $\mathcal{I}(\cdot)$ denotes a suitable interpolation operator.
In practice, we use a hybrid of linear and nearest-neighbour interpolation to preserve continuity in densely sampled regions while maintaining robustness in sparsely sampled areas. The 3D vertex coordinates $(x,y,z)$ for a mesh are projected directly to the 2D grid of shape $(3,50,50)$ where the first dimension (3) represents the channels corresponding to the $x$, $y$, and $z$ coordinates, and the latter two dimensions ($50\times 50$) define the resolution of the structured grid in the UAC system.
Fibre orientations, however, are originally defined per mesh element rather than per vertex. To obtain vertex-wise fibre directions suitable for UAC projection, we convert element-level vectors to nodal vectors using area-weighted averaging. For each vertex, we accumulate the fibre vectors from all incident triangles, scaling each vector by the area of its corresponding triangle so that larger elements contribute proportionally. The resulting vectors are normalised to unit length to preserve directional information. These vertex-level fibre directions are then mapped to the 2D grid using the same UAC interpolation procedure as these coordinates, resulting in a 2D grid of shape $(3,50,50)$.
4.1.3 Identify pacing locations
To ensure the generalisability of neural operators and to capture a wide range of conduction dynamics, the EP simulations were initiated using a distributed set of seven clinically relevant pacing sites in the LA: the Left Atrial Appendage (LAA), the Left Superior and Inferior Pulmonary Veins (LSPV, LIPV), the Right Superior and Inferior Pulmonary Veins (RSPV, RIPV), the Coronary Sinus (CS), and the Roof which is defined as the superior wall of the left atrium connecting the LSPV and RSPV [romero2017atrial, moreira2008atrial, elmariah2013coronary].
Pacing locations are defined across all patient anatomies using the UAC system. For each anatomical region, a 2D bounding box within the UAC domain was specified, delineating a subset of mesh vertices. The physical pacing centre is computed as the mean position of all vertices within that UAC-defined box. The stimulation region is subsequently defined by all mesh vertices located within a fixed Euclidean radius of $2.0\,\text{mm}$ from this central point, ensuring comparable stimulation volumes across all seven sites. The spatial distribution of these seven sites on a representative LA mesh is mapped onto the UAC coordinate system in Fig. 4.
Figure 4: Marking pacing sites on an LA. a, Anterior and posterior views of an LA mesh where the seven pacing locations have been marked. b, Spatial distribution of these pacing sites in UAC coordinates. The blue regions indicate the anatomical openings (pulmonary veins and mitral valve).
4.2 GPU-based FEM EP simulation
To generate a large-scale, comprehensive dataset for training, we simulated LA electrical propagation using the widely adopted monodomain equation [geselowitz1983bidomain] (further implementation details are provided in Supplementary Section 1). This biophysical model captures the electrical conduction through anisotropic cardiac tissue, balancing capacitive and ionic currents with diffusive current flow, governed by the longitudinal ($\sigma_{l}$) and transverse ($\sigma_{t}$) conductivities.
The primary and clinically significant output extracted from these simulations is the Local Activation Time (LAT) map, which defines the time $t$ at which the transmembrane potential $V$ first crosses a predefined threshold ($V_{\text{th}}$, typically 0 mV) at any location $x$ in the cardiac domain:
$\mathrm{LAT}(x)=\min\left\{t>0\ \middle|\ V(x,t)>V_{\text{th}}\right\},$
For each patient, we varied the longitudinal ($\sigma_{l}$) and transverse ($\sigma_{t}$) conductivities across seven clinically relevant pacing locations. At each pacing site, 300 combinations of $\sigma_{l}$ and the anisotropy ratio $\sigma_{l}/\sigma_{t}$ were sampled using Latin hypercube sampling [mckay2000comparison]. The chosen range of $\sigma_{l}$ (0.1–0.4 S/m) is consistent with experimentally measured conduction velocities and previously reported myocardial conductivity values [fu2024progress], while anisotropy ratios between 4 and 10 reflect the substantially faster fibre-aligned conduction observed in atrial tissue [kotadia2020anisotropic]. These values imply corresponding $\sigma_{t}$ in the range 0.01–0.1 S/m.
Each simulation was run for 600 ms to capture the full activation sequence, and the local activation time (LAT) map was recorded during runtime. Each simulation required approximately one minute on an NVIDIA A100 GPU, enabling the generation of 210,000 simulations for Cohort A and 98,700 simulations for Cohort B, yielding a combined total of 308,700 simulations. Executing these simulations across eight A100 GPUs reduced the total wall-clock time from an estimated 214 days (single GPU) to approximately 27 days.
4.3 Neural operator learning
We formulate LAT map prediction as a neural operator learning problem, in which the goal is to approximate a mapping from patient-specific anatomical and electrophysiological parameters to the corresponding spatial distribution of activation times. Formally, we aim to learn an operator
$\mathcal{G}:(A,\sigma,p)\mapsto\mathrm{LAT}(\mathbf{x}),$
where $A$ denotes the atrial geometry, $\sigma$ the conductivity tensor, $p$ the pacing site location, and $\mathrm{LAT}(\mathbf{x})$ the activation time at position $\mathbf{x}\in\Omega$ where $\Omega$ is the computational c. This framework requires the generation of large and diverse training datasets that pair input configurations $(A,\sigma,p)$ with ground-truth LAT maps obtained from high-fidelity EP simulations.
4.3.1 Model architecture
Our proposed architecture integrates a coordinate embedding module for point cloud projection with a Vision Transformer-based encoder and a convolutional decoder, enabling high-fidelity spatial field prediction from both sparse and structured physiological data. The model is designed to process heterogeneous inputs, including sparse point clouds (e.g., pacing coordinates) and spatially distributed scalar fields (e.g., tissue conductivity, surface area), by transforming them into a unified, fixed-resolution tensor representation suitable for transformer-based spatial modelling (Fig. 5).
Figure 5: Model architecture. a, The construction of the input and output for neural operator learning, where patient-specific anatomical and physiological parameters are projected onto a structured grid via the Universal Atrial Coordinate system. The pacing location is converted from a point cloud representation into a spatial distribution on the same grid, enabling consistent integration with other input modalities. b, A neural network module that maps a point cloud of shape $(N,2)$, representing spatial input features such as pacing sites or anatomical landmarks, into a structured 2D grid of shape $(1,50,50)$. Here, NN refers to the number of vertices where stimulus (pacing) is applied. c, The Vision Transformer block used in (d), which is the original ViT model proposed in [vit]. d, Our full encoder-decoder architecture, which uses convolutional patch embedding and a Vision Transformer encoder to process the input grid, and then upsamples the latent features using a transposed convolution and projects them into the final LAT map using a convolutional output layer.
Point-to-grid projection.
To encode pacing (stimulation) locations provided as sparse, unordered two-dimensional point clouds of shape $(N,2)$ where $N$ is the number of vertices in the pacing region, we introduce an embedding module that maps the point set into a 2D grid representation (Fig. 5b). This module follows the architectural design of PointNet [qi2017pointnet], applying a multi-layer perceptron (MLP) independently to each input point, followed by a permutation-invariant mean pooling operation to obtain a global descriptor of the stimulation pattern. This descriptor is then decoded via a learnable mapping to a 2D spatial grid of shape $(3,50,50)$. Unlike PointNet, which is optimised for classification and segmentation tasks, our adaptation is explicitly designed to produce spatially structured embeddings that are compatible with convolutional and transformer-based feature extractors.
Multi-modal feature integration.
The input to the neural operator is a multi-modal, multi-channel grid that encodes patient-specific mesh properties and simulation parameters into a fixed-size ten-channel representation (Fig. 5a), which can be divided into two groups. The first group comprises the seven geometric and anatomical channels, including the UAC-projected 3D vertex coordinates $(x,y,z)$ and the 3D fibre direction vectors (six channels in total, each forming a $(3,50,50)$ feature map). This is supplemented by the total surface area of the left atrium, a global anatomical feature that is broadcast across the spatial domain to form a $(1,50,50)$ constant grid. The second group comprises the two parametric features including the two conduction velocities ($\sigma_{l}$ and $\sigma_{t}$), which are also broadcast across the spatial domain to form a $(2,50,50)$ feature map, and the encoded pacing location in a $(1,50,50)$ feature map. The output of the neural network is trained against the target LAT map of shape $(1,50,50)$.
ViT-Based Spatial Representation Learning.
To model spatial dependencies across the input domain, we employ a Vision Transformer architecture that encodes the multi-channel input grid into a sequence of contextualised latent representations. A convolutional patch embedding layer [wu2021cvt] first divides the input grid into non-overlapping patches and projects each patch into a shared embedding space, producing a sequence of latent tokens. Learnable positional embeddings are added to retain spatial structure within the tokenised representation. This sequence is then processed by a stack of standard ViT blocks, each comprising multi-head self-attention and feed-forward sublayers, which enables the integration of global context and the capture of long-range spatial interactions (Fig. 5c).
Following the ViT-based encoder, the latent tokens are reshaped into a compact 2D feature map. To recover full spatial resolution, we employ a lightweight decoder consisting of a transposed convolution layer for upsampling, followed by a convolutional projection head that produces the final output map. The transposed convolution operation, originally introduced in deep learning by [zeiler2011adaptive] for adaptive deconvolutional networks and later popularised for CNN visualisation by [zeiler2014visualizing], enables learnable upsampling by applying the transpose of the convolution operator. This decoding step ensures that the global features encoded by the transformer are effectively translated back into a high-resolution spatial field for pixel-wise prediction.
4.3.2 Sobolev Loss with Spatial Smoothness Regularisations
To ensure our model produces physiologically consistent LAT maps, we employ an objective function grounded in the relative Sobolev $H^{1}$ norm. This choice enforces consistency not only in the predicted values but also in their spatial derivatives, a property essential for accurately modelling continuous and smooth conduction patterns.
To further refine the solution and mitigate non-physical, high-frequency artefacts while preserving sharp wavefront transitions, we incorporate two standard regularisation terms. The first is Total Variation (TV) regularisation $\mathcal{L}_{\mathrm{TV}}$, which penalises the $L^{1}$ norm of the image gradient. The second is a Laplacian regularisation $\mathcal{L}_{\mathrm{Laplacian}}$, which promotes harmonicity by penalising spatial curvature (full definitions are provided in Supplementary Section 4).
The overall objective function is defined as a weighted sum:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\mathrm{H}^{1}}(u,v)+\lambda\mathcal{L}_{\mathrm{TV}}(u)+\lambda\mathcal{L}_{\mathrm{Laplacian}}(u),$
where $\mathcal{L}_{\mathrm{H}^{1}}(u,v)$ is the primary loss between the predicted LAT map ($u$) and the ground truth ($v$), and the hyperparameter $\lambda$ controls the strength of the regularisation. This combination ensures that the model’s predictions are both highly accurate and physiologically coherent.
Declarations
Funding
This work is supported by the NIHR Imperial Biomedical Research Centre (BRC) and by the British Heart Foundation Centre of Research Excellence (RE/24/130023). SAN is supported by NIH R01-HL152256 and R01-HL162260, ERC PREDICT-HF453 (864055), BHF (RG/20/4/34803), EPSRC (EP/P01268X/1, EP/Z531297/1) and by the Technology Missions Fund under the EPSRC Grant EP/X03870X/1 & The Alan Turing Institute. CR receives funding from the British Heart Foundation (RG/20/4/34803). We acknowledge computational resources and support provided by the Imperial College Research Computing Service.
Data availability
Dataset A used in this study is publicly available on Figshare under the permanent identifier 10.6084/m9.figshare.30712559. Dataset B contains sensitive information and cannot be publicly released due to ethical and privacy restrictions. Access to Dataset B may be requested from the corresponding author and will be provided subject to institutional approval and a formal data-sharing agreement.
Code availability
All code used to generate the results in this study is available at: https://github.com/sagebei/unified_framework_operator_learning.

### v2

@@
 Abstract
-Accurate maps of atrial electrical activation are essential for personalised treatment of arrhythmias, yet biophysically detailed simulations remain computationally intensive for real-time clinical use or population-scale analyses.
-Here we introduce a geometry-independent operator-learning framework that predicts local activation time (LAT) fields across diverse left atrial anatomies with near-instantaneous inference.
-We generated a dataset of 308,700 simulations using a GPU-accelerated electrophysiology solver, systematically varying multiple pacing sites and physiologically varied conduction properties across 147 patient-specific geometries derived from two independent clinical cohorts.
-All anatomical and functional data are expressed in a Universal Atrium Coordinate system, providing a consistent representation that decouples electrophysiological patterns from mesh topology.
-Within this coordinate space, we designed a neural operator with a vision-transformer backbone to learn the mapping from structural and electrophysiological inputs to LAT fields.
-With a mean prediction error of 5.1 ms over a 455 ms maximum simulation time, the model outperforms established operator-learning approaches and performs inference in 0.12 ms per sample.
-Our framework establishes a general strategy for learning domain-invariant biophysical mappings across variable anatomical domains and enables integration of computational electrophysiology into real-time and large-scale clinical workflows.
+Learning biophysically accurate solution operators for cardiac electrophysiology is fundamentally challenged by geometric variability across patient-specific heart anatomies.
+Most existing neural operator approaches are limited to structured or weakly deformed domains, restricting their applicability to realistic atrial and ventricular geometries.
+Here, we introduce a unified operator-learning framework that projects inputs and outputs onto a standardised anatomical coordinate system, decoupling electrophysiological dynamics from mesh topology.
+This formulation enables geometry-independent learning while preserving physiologically meaningful spatial organisation, and allows predictions to be interpolated back onto patient-specific geometries for anatomical interpretation.
+To support large-scale training within the framework, we develop a GPU-accelerated electrophysiology solver and generate over 300,000 high-fidelity simulations across diverse patient-specific left atrial geometries with varied pacing and conduction properties.
+Within this anatomical coordinate domain, we design a neural operator to predict full-field local activation time maps, achieving a mean absolute error of 5.1 ms and an inference time of 0.12 ms per sample, outperforming existing operator learning and convolutional baselines.
+We further validate the framework on ventricular geometries, demonstrating robust generalisation beyond the atrial setting.
+Together, this framework establishes a scalable foundation for fast, geometry-invariant cardiac electrophysiology modelling, with potential relevance for real-time and population-scale clinical workflows.
 1 Introduction
-Atrial fibrillation (AF) is the most common sustained cardiac arrhythmia, affecting around 46.3 million individuals worldwide [kornej2020epidemiology], and represents a major cause of stroke [cheng2024global], heart failure [li2025global], and sudden cardiac death [lai2025causes].
-Computational simulations have emerged as powerful tools in AF research and clinical trials, offering insights into arrhythmia mechanisms, enabling the prediction of ablation outcomes, and personalised therapeutic strategies [heijman2021computational, boyle2019computationally, roney2021time, krogh2012effects].
-Many of these applications rely on calibrating patient-specific biophysical models to local activation time (LAT) maps, which capture the spatiotemporal propagation of electrical wavefronts across the myocardium.
-This calibration typically involves solving large ensembles of high-fidelity finite element method (FEM) simulations that numerically solve the governing partial differential equations (PDEs) of cardiac electrophysiology on anatomically accurate patient-specific geometries [goktepe2009computational, rocha2010macro].
-Although highly accurate, FEM workflows are computationally expensive and often take hours to days to solve, making them impractical for intraoperative use.
-Consequently, patient-specific models are usually prepared in advance from pre-procedure MRI or CT scans that define anatomy and tissue properties.
-In many clinical scenarios, however, such imaging is unavailable, and the only anatomical and electrophysiological data come from intraoperative electro-anatomical mapping (EAM).
-While EAM-derived LATs can be used for retrospective calibration [jaffery2024review], the turnaround time of FEM-based methods prevents their use for real-time decision-making during ablation.
-Recent advances in machine learning have opened new opportunities to accelerate the numerical solution of cardiac electrophysiology models.
-Physics-informed neural networks (PINNs) have been proposed as a physics-constrained alternative to purely data-driven surrogates, embedding the governing PDE residuals directly into the loss function [pinn].
-They are typically trained for a single PDE configuration, defined by a fixed geometry, parameter set, and boundary conditions [kovachki2023neural].
-However, they do not naturally generalise to new configurations without retraining: if the domain shape, material parameters, or boundary conditions change, the PINN must be re-trained from scratch or at least fine-tuned.
-Neural operators directly address this limitation by learning a mapping between infinite-dimensional function spaces rather than a single instance of a PDE [kovachki2023neural].
-Foundational neural operators such as the Deep Operator Network (DeepONet) [don] and the Fourier Neural Operator (FNO) [fno] have demonstrated strong performance on canonical problems defined over structured domains.
-By learning this operator mapping, they can efficiently predict solution fields for new inputs, even for unseen geometries or different discretisation resolutions, without the need for time-consuming retraining.
-Applying neural operators to complex real-world problems, such as predicting cardiac activation maps, presents significant challenges.
-First, these aforementioned neural operator frameworks are built to handle data on structured, canonical domains, making them ill-suited for diverse and irregular geometries [geofno, gino].
-Second, training a reliable neural operator is contingent on the availability of large, high-quality datasets of diverse simulation results [ahmad2024diffeomorphic], and generating such datasets is a major challenge [trayanova2021machine].
-Several approaches have been proposed to address the challenge of anatomical variability.
-Diffeomorphic Mapping Operator Learning (DIMON) [yin2024scalable] uses diffeomorphic transformations to map individual geometries to a shared, standardised space.
-This allows a neural operator to learn latent solutions on the reference domain, with the original shape encoded as an auxiliary input.
-While effective for relatively simple geometries, such as the semi-spherical left ventricle (LV), this strategy faces substantial difficulties when extended to anatomies with more intricate topologies, such as the multi-holed structure of the left atrium, where constructing topologically consistent reference domains is non-trivial.
-Similarly, methods like GeoFNO require a diffeomorphic mapping to simple domains (e.g., cubes or tori) where the Fourier transform can be applied [geofno]; however, the topologically complex 3D geometries of the human left atrium, with multiple openings, make such mappings impossible [loeffler2024graph].
-Graph-based methods like the Graph Fourier Neural Kernel (G-FuNK) [loeffler2024graph] learn directly on mesh graphs, avoiding domain mapping.
-However, their spectral bases are sensitive to mesh quality and topology, reducing robustness and generalisability in heterogeneous clinical datasets.
-Here, we introduce a method that overcomes the dependency of neural operators on structured domains for electrophysiology simulations.
-Our approach projects inputs and outputs onto a common, anatomically standardised reference using the Universal Atrial Coordinates (UAC) system [uac].
-The UAC defines a mesh-independent coordinate space for the left atrium (LA), preserving physiologically relevant spatial relationships while unifying geometry and topology across patients.
-This representation enables neural operators such as DeepONet and FNO to consistently interpret spatial features, free from irregular mesh geometries or topological inconsistencies.
-To support large-scale training, we developed a GPU-accelerated FEM solver, drastically reducing the time required to compute LAT maps and enabling the generation of a high-fidelity, comprehensive dataset.
-Using this dataset, we train a neural operator to approximate the underlying biophysical mapping, achieving near real-time inference without compromising physiological accuracy.
-Our model, an encoder-decoder architecture with a Vision Transformer (ViT) [vit] backbone, predicts full-field LAT maps from patient-specific anatomical features, conductivity, and pacing sites.
-It achieves a mean absolute error of $5.1\pm 0.2$ ms relative to FEM simulations, exhibiting high pointwise accuracy, spatial smoothness, and robust generalisation across independent clinical cohorts, overcoming the domain shift inherent in multi-centre medical data.
-By integrating GPU-accelerated simulation and neural operator training into a unified framework, our approach addresses both the computational bottleneck of large-scale electrophysiology simulations and the geometric variability of patient-specific atrial anatomies.
-It enables accurate prediction of full-field LAT maps in just 0.12 ms per sample on a GPU, offering a pathway toward rapid, patient-specific, physiology-informed decision support for AF interventions.
+Computational models of cardiac electrophysiology provide a principled framework for studying electrical wave propagation, investigating arrhythmia mechanisms, and informing patient-specific therapeutic interventions [niederer2019computational, trayanova2024computational, prakosa2018personalized].
+A central application of these models is the analysis of local activation time (LAT) maps, which characterise the spatiotemporal evolution of electrical activation across the myocardium [williams2018local, jaffery2024review, karoui2021cardiac].
+Generating and analysing such LAT maps typically requires large ensembles of high-fidelity finite element method (FEM) simulations on anatomically detailed, patient-specific geometries, involving the numerical solution of coupled partial differential equations governing cardiac electrophysiology [niederer2011verification, opencarp].
+Atrial fibrillation (AF) is a primary clinical setting in which such patient-specific LAT analyses are routinely required [sahli2018interpreting].
+As the most common sustained cardiac arrhythmia, AF affects approximately 46.3 million individuals worldwide [kornej2020epidemiology] and is a leading cause of stroke [cheng2024global], heart failure [li2025global], and sudden cardiac death [lai2025causes].
+Computational modelling has therefore become an important tool in AF research and clinical studies, supporting mechanistic investigation of arrhythmia drivers, evaluation of ablation strategies, and development of personalised treatment planning [heijman2021computational, boyle2019computationally, roney2021time, krogh2012effects].
+However, the substantial computational cost of FEM-based workflows, often requiring hours to days per simulation, limits their applicability in settings that demand rapid or real-time decision-making [niederer2011simulating, boyle2019computationally].
+Recent advances in machine learning have opened new avenues for accelerating cardiac electrophysiology simulations.
+Neural operators, in particular, aim to learn mappings between infinite-dimensional function spaces [kovachki2023neural], enabling solution fields to be predicted for unseen inputs, discretisations, or geometries without retraining.
+Architectures such as DeepONet [don] and the Fourier Neural Operator (FNO) [fno] have demonstrated strong performance on parametric partial differential equation problems.
+Despite this promise, most existing neural operator frameworks are developed for structured or weakly deformed computational domains, which limits their applicability to the highly irregular, patient-specific geometries encountered in cardiac modelling [geofno, gino].
+In addition, training robust and generalisable operators requires large, high-quality simulation datasets, which remain expensive to generate using conventional electrophysiology solvers [ahmad2024diffeomorphic, trayanova2021machine].
+Several strategies have been proposed to address anatomical variability in operator learning.
+Diffeomorphic mapping approaches [zhao2025diffeomorphism], including Diffeomorphic Mapping Operator Learning (DIMON) [yin2024scalable], map individual geometries to a shared reference domain, enabling learning in a common representation while encoding geometry as auxiliary information.
+While effective for relatively simple geometries, such as the semi-spherical left ventricle, these approaches become challenging for anatomies with complex topology, such as the multi-holed left atrium, where constructing globally consistent mappings is non-trivial.
+Related methods, such as Geo-FNO [geofno], deform irregular geometries to structured domains to enable spectral learning, but the accuracy and stability of such mappings may deteriorate for highly intricate anatomical structures.
+Here, we propose a framework that reformulates neural operator learning in a coordinate domain intrinsic to cardiac anatomy rather than in the space of mesh discretisations.
+By expressing model inputs and outputs in a shared anatomical parameterisation, electrophysiological patterns can be represented consistently across patients despite substantial geometric and topological variability.
+Predictions obtained in this coordinate space can subsequently be mapped back onto the original patient-specific geometry, enabling anatomically faithful reconstruction and direct clinical interpretation.
+In this work, we adopt the Universal Atrial Coordinates (UAC) [uac] and Universal Ventricular Coordinates (UVC) [uvc] systems to define such intrinsic representations for atrial and ventricular geometries, respectively.
+These coordinate systems preserve physiologically meaningful spatial organisation while providing a unified domain across diverse anatomies, enabling neural operators to learn spatially coherent solution mappings without sensitivity to irregular mesh connectivity or topological inconsistencies.
+To support operator training within this framework, we further develop a GPU-accelerated electrophysiology simulation pipeline that enables efficient generation of high-fidelity LAT maps at scale across diverse patient-specific geometries.
+By integrating high-performance simulation with neural operator learning in anatomical coordinate space, our approach addresses both the computational cost of cardiac electrophysiology modelling and the geometric variability inherent in patient-specific anatomies, providing a scalable pathway toward rapid, physiology-informed inference in atrial fibrillation.
 2 Results
@@
 This domain shift has a measurable impact on cross-domain generalisation performance.
+The methodology used to assess this effect is detailed in Methods.
 We next quantified the primary determinants of total activation time across all simulations.
@@
 The notation $X\rightarrow Y$ indicates that the model was trained on source dataset $X$ and evaluated on target dataset $Y$.
-We evaluated the contribution of each anatomical input modality by removing spatial coordinates, fibre orientation, and surface area features in turn, and retraining the model under each condition.
-Performance was assessed using two complementary metrics: Mean Absolute Error (MAE), which quantifies the numerical accuracy (error in milliseconds), and the Structural Similarity Index Measure (SSIM) [wang2004image], which measures the spatial fidelity and structural quality of the predicted activation maps.
-All analyses in this section were performed using 5-fold cross-validation.
-Unless stated otherwise, experiments were conducted on Dataset A, with cross-domain evaluations involving both Datasets A and B.
-The results in Fig.
-2a show that incorporating all three feature types achieves accurate and spatially smooth LAT predictions (MAE $6.09\pm 0.26$ ms; SSIM $0.963\pm 0.001$).
-Removing the surface area feature results in a pronounced degradation in performance (MAE $9.88\pm 0.81$ ms), indicating that global anatomical scaling is an essential contextual signal for correct wavefront organisation.
-Similarly, excluding spatial coordinates reduces both numerical accuracy and structural similarity (MAE $9.04\pm 0.44$ ms; SSIM $0.948\pm 0.002$), demonstrating that explicit localisation in the UAC domain is critical for resolving conduction pathways.
-In contrast, removing fibre orientation led to improved performance across both metrics (MAE $5.15\pm 0.25$ ms; SSIM $0.978\pm 0.002$).
-This suggests that the fibre orientation information, as currently represented, may introduce noise or redundancy relative to the geometric cues already encoded in the coordinate and surface area channels.
-We discuss this observation further in the discussion section.
-Accordingly, fibre orientation features were excluded from all subsequent experiments.
+We quantified the contribution of individual anatomical input modalities by systematically removing spatial coordinates, fibre orientation, and surface area features, retraining the model under each configuration.
+Model performance was evaluated using mean absolute error (MAE), which captures numerical accuracy in milliseconds, and the structural similarity index measure (SSIM) [wang2004image], which assesses spatial fidelity of the predicted activation maps.
+All results are reported using five-fold cross-validation.
+Unless otherwise stated, experiments were conducted on Dataset A, with cross-domain evaluations involving both Datasets A and B.
+As shown in Fig.
+2a, incorporating all anatomical features yields accurate and spatially coherent predictions of local activation time.
+Removing the surface area feature leads to a substantial degradation in performance, highlighting the importance of global anatomical scaling for correctly organising activation wavefronts.
+Similarly, excluding spatial coordinates significantly reduces both numerical accuracy and structural similarity, demonstrating that explicit localisation in the UAC domain is essential for resolving conduction pathways.
+In contrast, removing fibre orientation features unexpectedly improves predictive performance across both metrics.
+This observation indicates that, in the present formulation, fibre orientation does not provide an additional informative signal beyond that already captured by geometric coordinates and global anatomical features.
+Based on this result, fibre orientation inputs were excluded from all subsequent experiments.
+Explanations for this effect are discussed further below.
 2.3 Evaluating the effect of regularisation
-To assess the impact of regularisations on LAT map reconstruction, we evaluated the contribution of two common spatial regularisation terms, total variation and second-order Laplacian loss, when added individually or in combination.
-These regularizers aim to encourage spatial smoothness and suppress high-frequency artefacts in the predicted LAT maps.
-To evaluate the impact of different loss formulations on model performance, we conducted systematic experiments by varying the regularisation coefficient $\lambda\in\{5\times 10^{-2},3\times 10^{-2},1\times 10^{-2},5\times 10^{-3},1\times 10^{-3}\}$ in the loss function.
-For each fold in 5-fold cross-validation, the model achieving the best performance on the validation set was retained, and the corresponding metrics were recorded.
-The performance for a given regularisation weight $\lambda$ was derived by averaging these metrics across the five folds.
-To probe the effect of each regularisation configuration, such as TV alone or the combination of TV and Laplacian, we further averaged the fold-averaged metrics across all tested $\lambda$ values (Fig.
+We investigated the impact of spatial regularisation on LAT map reconstruction by augmenting the $H^{1}$ loss with two commonly used smoothness terms: total variation (TV) and a second-order Laplacian penalty, applied individually or in combination.
+These regularisation terms are designed to suppress high-frequency artefacts while preserving physiologically meaningful spatial structure.
+Model performance was evaluated across a range of regularisation strengths ($\lambda\in\{5\times 10^{-2},3\times 10^{-2},1\times 10^{-2},5\times 10^{-3},1\times 10^{-3}\}$) using five-fold cross-validation.
+For each configuration, performance metrics were averaged across folds and regularisation weights to isolate the effect of the regularisation formulation (Fig.
 2b).
-Among all evaluated configurations, the $H^{1}$ loss combined with both TV and Laplacian regularisations achieved the best overall performance, with a mean MAE of $5.16$ ms and a mean SSIM of $0.978$.
-This demonstrates that combining both regularisation terms effectively enhances structural fidelity while preserving numerical accuracy.
-The addition of TV or Laplacian individually yielded smaller improvements, suggesting that their combined effect is necessary to fully capture the relevant spatial structures.
+Among all tested configurations, combining the $H^{1}$ loss with both TV and Laplacian regularisation consistently yielded the best performance, achieving the lowest mean MAE and highest SSIM.
+This result indicates that jointly enforcing first- and second-order smoothness constraints enhances structural fidelity without sacrificing numerical accuracy.
+In contrast, applying either TV or Laplacian regularisation alone resulted in smaller and less consistent improvements, suggesting that their combined effect is necessary to adequately capture the spatial organisation of activation patterns.
 2.4 Benchmarking model performance
-We benchmarked a range of neural operator frameworks and deep-learning architectures capable of modelling cardiac activation patterns from multimodal inputs.
-For each model, we selected three primary hyperparameters governing capacity and constructed six configurations by varying these across representative ranges, ensuring that the comparison is both systematic and approximately capacity-matched across architectures (Fig.
+We benchmarked our approach against a range of neural operators and convolutional architectures commonly used for modelling cardiac activation from multimodal inputs.
+To ensure a fair comparison, models were evaluated across approximately capacity-matched configurations spanning representative hyperparameter ranges (Fig.
 2c).
-Supplementary Section 2 provides further details on architecture design and training procedures.
-We first evaluated operator learning approaches.
-While the Fourier Neural Operator (FNO) [fno] is theoretically well-suited for learning global mappings in function space, its empirical performance was moderate (MAE $10.14\pm 0.07$ and SSIM $0.938\pm 0.005$), indicating potential challenges in handling heterogeneous and noisy physiological data.
-The Wavelet Neural Operator (WNO) [wno] exhibited the weakest performance among evaluated models (MAE $10.08\pm 0.28$ and SSIM $0.938\pm 0.007$, likely reflecting sensitivity to measurement noise and anatomical variability inherent in real-world datasets.
-The DeepONet framework, using our proposed ViT-based architecture for both branch and trunk networks, achieved strong performance (MAE $7.12\pm 0.37$ ms; SSIM $0.963\pm 0.005$).
-Among convolutional baselines, the U-Net architecture [unet] outperformed the simpler ResNet model [resnet].
-The U-Net’s multi-scale encoder-decoder structure, with its strong inductive bias for spatial localisation, yielded a balanced performance of MAE $8.48\pm 0.28$ ms and SSIM $0.965\pm 0.004$.
-However, these convolutional models were fundamentally limited in their ability to capture global pacing effects and long-range spatial dependencies
-Our model surpassed all baselines, achieving the best results with an MAE of $5.15\pm 0.25\text{ ms}$ and SSIM of $0.978\pm 0.002$.
-This confirms the model’s superior capability to synthesise heterogeneous inputs into spatially coherent, smooth, and highly accurate LAT map predictions.
-While MAE and SSIM provide global measures of model performance, they do not reveal where errors concentrate spatially.
-We therefore visualised the prediction error distributions across the UAC domain.
-Fig.
-2d displays the spatial distribution of prediction error on LAT maps for six test cases, each averaged over seven pacing locations and 300 combinations of conductivity velocities.
-Across all six test cases, a visually consistent and highly coherent source of localised prediction error is observed near the horizontal central line of the UAC map.
-This localised band of elevated error is the most prominent spatial feature in the error distribution plots.
-Beyond this central band, the prediction errors generally remain low and diffuse across the rest of the UAC domain, indicating high model fidelity in most tissue regions.
-Beyond predictive accuracy, we further assessed the computational efficiency of our model to evaluate its suitability for real-time applications.
-Inference on batches of 1000 samples was performed on an NVIDIA RTX 5090 GPU in evaluation mode with CUDA synchronisation to ensure precise wall-clock timing.
-The model achieved an average inference time of $122.23$ ms per batch, or approximately $0.12$ ms per sample.
-This represents a five to six orders of magnitude acceleration compared with conventional finite-element electrophysiology solvers, which require approximately one minute per simulation with our solver in this study.
-Such efficiency enables near real-time prediction of LAT maps across diverse atrial anatomies and pacing configurations, highlighting the potential of our model for integration into time-critical clinical workflows and large-scale population studies.
+Full architectural details and exhaustive results are provided in Supplementary Section 2 and Supplementary Table 1.
+Among operator learning methods, the Fourier Neural Operator (FNO) and Wavelet Neural Operator (WNO) exhibited limited empirical performance in this setting, indicating challenges in capturing heterogeneous and noisy electrophysiological patterns on complex anatomical domains.
+DeepONet, implemented using our ViT-based backbone for both branch and trunk networks, achieved substantially improved accuracy, reflecting its stronger capacity to model global functional relationships.
+Convolutional baselines showed a similar trend.
+The U-Net architecture outperformed the simpler ResNet model, consistent with its multi-scale encoder–decoder structure and strong inductive bias for spatial localisation.
+Nevertheless, convolutional models remained limited in their ability to capture global pacing effects and long-range spatial dependencies.
+Our model outperformed all baselines, achieving the highest numerical accuracy and structural fidelity across all evaluated metrics.
+This result demonstrates its superior ability to integrate heterogeneous anatomical and physiological inputs into spatially coherent and physiologically consistent LAT predictions.
+To assess where residual errors concentrate spatially, we visualised prediction error distributions across the UAC domain (Fig.
+2d).
+Across representative test cases, errors are predominantly localised to a narrow horizontal band corresponding to the UAC seam, while remaining low and diffuse elsewhere, indicating high fidelity across most anatomical regions.
+Beyond accuracy, we evaluated computational efficiency to assess suitability for real-time applications.
+On a modern NVIDIA GPU, the model achieved an average inference time of approximately $0.12$ ms per sample.
+This represents orders of magnitude acceleration relative to conventional finite-element electrophysiology solvers, enabling near real-time LAT prediction across diverse anatomies and pacing configurations.
 2.5 Model generalisation across pacing sites
-Understanding whether a single model can generalise across diverse pacing conditions is critical for building scalable and clinically useful predictors of atrial activation, as it allows us to determine whether it can match or exceed the performance of individually trained models.
-Success in this setting would indicate that the model captures shared electrophysiological structure, enabling robust predictions without prior knowledge of stimulation location.
+Generalisation across diverse pacing conditions is critical for scalable and clinically useful prediction of atrial activation, as it determines whether a single model can match or exceed the performance of models trained for individual pacing sites.
+A model that performs well across pacing sites indicates that it has learned shared electrophysiological structure rather than site-specific patterns.
 We thus assessed the predictive performance of two modelling paradigms for LAT map reconstruction: (i) Single Pacing Location Models, trained independently for each of seven pacing sites, and (ii) a Unified Model, trained jointly across all sites to learn a shared inductive bias over anatomical and physiological structure.
@@
 Each evaluation was performed on models trained with 18 combinations of hyperparameters to ensure statistical significance, with results highlighting consistent trends (Fig.
-2f; full results in Supplementary Section 3).
+2f; full results are provided in Supplementary Tables 2 and 3).
 The strong internal performance confirms that the UAC projection effectively separates anatomical representation from patient-specific shape.
@@
 Training on the combined datasets $(A+B)$ encourages the model to focus on consistent anatomical features, ignoring spurious correlations tied to individual acquisition or annotation pipelines.
-Across both datasets, MAE decreased and SSIM increased in comparison with the internal performance, demonstrating that exposure to independent cohorts acts as a form of regularisation.
+Across both datasets, MAE decreased, and SSIM increased in comparison with the internal performance, demonstrating that exposure to independent cohorts acts as a form of regularisation.
 For completeness, we also evaluated the model trained and tested on the combined datasets $(A+B)\rightarrow(A+B)$, obtaining $(5.77\pm 0.16)$ for MAE and $(0.9734\pm 0.003)$ for SSIM.
+2.7 Interpolation to patient-specific geometry
+To assess anatomical prediction fidelity after interpolation, we evaluated model performance on both LA surface meshes and left ventricular (LV) volume meshes (Fig.
+3).
+Predictions produced in the intrinsic anatomical coordinate domain were interpolated back onto the original patient-specific mesh geometry, enabling anatomically faithful reconstruction in both settings.
+Figure 3: Interpolation of neural operator predictions to patient-specific atrial and ventricular mesh geometries.
+a, Left: local activation time (LAT) predictions interpolated from the coordinate domain back onto representative left atrial (LA) surface meshes.
+Regions where the absolute prediction error exceeds 15 ms are highlighted.
+Right: distribution of mean absolute error across pacing locations evaluated on 100 LA test cases.
+b, Corresponding results for left ventricular (LV) meshes, showing interpolated LAT predictions, regions of large absolute error, and error distributions across pacing locations.
+Model performance was evaluated using a previously proposed benchmark task in which LAT is predicted from one of seven pacing locations alone [yin2024scalable].
+Our experiments were conducted on Cohort A using 100 left atrial geometries, of which 20 were held out for testing.
+In addition, 1,006 left ventricular (LV) meshes were obtained from a publicly available dataset [lvdata], with 106 meshes reserved for evaluation.
+Model inputs and outputs were represented on a $100\times 100$ grid for LAs and a $50\times 50\times 50$ volumetric grid for LVs.
+On test data, the model achieves mean absolute errors of 5–6 ms for LA geometries, relative to maximum activation times of up to 285 ms.
+For LV geometries, errors are consistently lower, typically ranging between 2 and 4 ms for maximum activation times of up to 350 ms.
+These ventricular error levels improve upon those reported in prior geometry-independent operator learning approaches, where mean absolute errors of approximately 5–6 ms were observed for LV meshes [yin2024scalable].
+In both atrial and ventricular settings, errors remain consistent across pacing locations, indicating stable performance independent of stimulation site.
+Spatially, elevated errors in the LA are primarily localised to the atrial appendage and pulmonary vein openings, regions characterised by complex geometry and boundary effects.
+For the LV, larger errors are observed near the basal openings, corresponding to truncation at the valve plane, and along the interface with the right ventricle, reflecting geometric discontinuities introduced during extraction from bi-ventricular meshes.
 3 Discussion
-The increasing complexity of modern biomedical simulations, particularly in fields like computational cardiology [brown2025cardiac], has created a growing need for methods that can bridge the gap between high-fidelity mechanistic models and the demands of real-time clinical applications.
-The results demonstrate that our proposed framework can accurately and robustly learn the complex, nonlinear mapping from a patient’s anatomy and electrophysiological parameters to their full-field LAT map, achieving high fidelity to biophysical simulations in milliseconds.
-This represents an advancement over conventional approaches and a crucial step toward enabling in-procedure, real-time, personalised cardiac electrophysiology analysis.
 3.1 Geometry-independent neural operator learning
-Our approach demonstrates that a well-designed coordinate system is a powerful tool for bridging the gap between irregular, real-world data and the structured-domain requirements of neural networks, allowing a single neural operator to learn a geometry-independent mapping.
-By leveraging the UAC system, we established a canonical representation that enabled neural operators to robustly capture the complex, nonlinear physics of cardiac conduction [hao2023gnot, li2023transformer, shih2025transformers, ovadia2024vito].
-A key finding is the model’s low MAE of $(5.15\pm 0.25)$ ms.
-This level of accuracy is significant in a clinical context, as it is well within the typical measurement noise and physiological variability observed in cardiac electroanatomic mapping systems [gaeta2021high] (typically on the order of 5–10 ms [narayan2024advanced]).
-This demonstrates that our model’s predictions are not only numerically accurate but also fall within a clinically acceptable margin, making the approach viable for non-invasive patient stratification, procedure planning, and guiding ablation therapies.
-A central tenet of our work is the understanding that the success of a neural operator relies on the availability of a large, diverse, and high-quality training dataset.
-By developing a GPU-accelerated FEM solver, we were able to generate a large simulation dataset across diverse patient-specific left atrial anatomies, providing a robust foundation for learning generalizable principles of cardiac electrophysiology.
-This data generation pipeline is a critical component of our framework, as it bypasses the prohibitive computational bottleneck that has historically limited the application of data-driven methods in this domain.
+Our approach demonstrates that a well-designed coordinate system is an efficient tool for bridging the gap between irregular, real-world data and the structured-domain requirements of neural networks, allowing a single neural operator to learn a geometry-independent mapping.
+By leveraging the universal coordinate system, we established a canonical representation that enabled neural operators to robustly capture the complex, nonlinear physics of cardiac conduction [hao2023gnot, li2023transformer, shih2025transformers, ovadia2024vito].
+The resulting prediction accuracy, with mean absolute errors on the order of 5 ms, lies within the typical measurement uncertainty and physiological variability reported for clinical electroanatomic mapping systems [gaeta2021high] (typically on the order of 5–10 ms [narayan2024advanced]).
+This demonstrates that our model’s predictions are not only numerically accurate but also fall within a clinically acceptable margin.
+A central premise of this work is that effective neural operator learning in cardiac electrophysiology depends on access to large, diverse, and high-quality training data.
+To this end, we developed a GPU-accelerated finite element electrophysiology solver that enables efficient generation of high-fidelity simulations across a wide range of patient-specific left atrial anatomies and electrophysiological conditions.
+This data generation pipeline addresses a longstanding computational bottleneck in cardiac modelling and provides a scalable foundation for learning generalisable electrophysiological structure.
 Beyond supporting our framework, this large, curated dataset provides a foundation for future data-driven studies in cardiac electrophysiology.
 Its scale, diversity, and consistency enable predictive models to generalise across anatomies and protocols, a capability previously limited by data scarcity.
-For comparison, G-FuNK [loeffler2024graph], trained on 25 left atrial meshes from atrial fibrillation patients, concluded that additional geometries could improve performance.
-DIMON [yin2024scalable] used 1,006 left ventricular patient data, and each was simulated at seven pacing sites, totalling 7,042 simulations.
+For context, G-FuNK [loeffler2024graph] was trained on 25 left atrial geometries, while DIMON [yin2024scalable] leveraged 1,006 ventricular geometries simulated at multiple pacing sites.
 In contrast, our simulations preserve these locations while also varying tissue conductivity, allowing neural operators to capture electrophysiological responses across diverse stimulation patterns and heterogeneous, patient-specific left atrial anatomies.
 Overall, our dataset spans 308,700 simulations, offering a far richer foundation for training and benchmarking state-of-the-art neural operators.
+The observation that removing fibre orientation features improved model performance was initially counterintuitive.
+Our analysis suggests two primary reasons for this finding.
+First, the current representation of fibre orientation may contain inherent noise or inaccuracies that interfere with the model’s ability to learn the true underlying electrophysiological dynamics.
+Second, the model appears to be leveraging the spatial and geometric priors provided by the coordinates and surface area features more effectively.
+These geometric inputs may implicitly contain sufficient information about the overall anatomical structure, making the explicit fibre orientation data redundant or even counterproductive in its current form.
 3.2 Comparative model analysis
@@
 This finding establishes a generalizable strategy for achieving domain-invariant medical image analysis.
-Our UAC projection provides the necessary geometry-invariant foundation, while training on heterogeneous data acts as a regularization method in feature space, compelling the model to learn only the biophysically relevant structures common to both domains and emphasizing that true translational power derives from integrating diverse, multi-centre data rather than from isolated single-site optimisation.
+The universal coordinate projection provides the necessary geometry-invariant foundation, while training on heterogeneous data acts as a regularisation method in feature space, compelling the model to learn only the biophysically relevant structures common to both domains and emphasising that true translational power derives from integrating diverse, multi-centre data rather than from isolated single-site optimisation.
 3.4 Limitations and future directions
-Despite its strong performance, the present study has limitations.
-The observation that removing fibre orientation features improved model performance was initially counterintuitive.
-Our analysis suggests two primary reasons for this finding.
-First, the current representation of fibre orientation may contain inherent noise or inaccuracies that interfere with the model’s ability to learn the true underlying electrophysiological dynamics.
-Second, the model appears to be leveraging the spatial and geometric priors provided by the coordinates and surface area features more effectively.
-These geometric inputs may implicitly contain sufficient information about the overall anatomical structure, making the explicit fibre orientation data redundant or even counterproductive in its current form.
-Beyond the feature representation, limitations also arise from the use of the UAC system.
-The UAC system is designed to map the closed, complex 3D left atrial surface onto a flat, 2D grid.
-To achieve this flattening, a continuous line must be cut on the 3D surface.
-Depending on the chosen coordinate orientation, this seam manifests as a central line on the 2D map.
-This seam creates a spatial discontinuity for the learning model, as points that are physically adjacent on the 3D surface are severed across this line in the 2D domain.
-Consequently, the neural operator struggles to accurately interpolate the LAT function across this boundary, resulting in a consistent band of large error centred precisely along this necessary central UAC seam.
-The additional large localised errors, which are more distributed, are plausibly linked to areas of extreme anatomical complexity and steep electrophysiological gradients.
-These clusters primarily occur around the highly patient-specific structures like the PV junctions and LAA.
-The UAC projection cannot fully remove the residual geometric variability present in these complex regions, increasing the model’s uncertainty.
-Furthermore, activation wavefronts often exhibit the steepest LAT gradients in these areas.
-The model is challenged to precisely localise and reproduce these sharp functional variations across diverse anatomies, leading to the small, perceptible clusters of higher error in these critical physiological locations.
-These difficulties suggest that while our model excels at capturing global activation patterns, the fine-grained, localised details influenced by complex anatomical geometries remain a challenging frontier.
-Future work could explore incorporating a more fine-grained representation of these critical regions to further enhance local accuracy.
+Beyond feature representation, limitations also arise from the use of intrinsic anatomical coordinate systems.
+For the left atrium, the UAC system maps a closed, topologically complex three-dimensional surface onto a flat two-dimensional domain.
+This flattening necessarily requires introducing a cut line on the surface, which appears as a seam in the coordinate representation.
+As a result, points that are physically adjacent on the original anatomy become separated across this seam, creating an artificial spatial discontinuity for the learning model.
+Consistent with this construction, our results reveal a narrow, coherent band of elevated prediction error aligned with the UAC seam, indicating that the neural operator struggles to interpolate smoothly across this boundary.
+In addition to seam-related effects, further localised errors are observed in regions of pronounced anatomical and electrophysiological complexity, particularly around the pulmonary vein junctions and the left atrial appendage.
+These regions exhibit substantial inter-patient variability that cannot be fully normalised by the UAC projection, leading to residual geometric heterogeneity in the coordinate domain.
+Moreover, activation wavefronts in these areas often display steep local gradients in activation time, increasing sensitivity to small spatial misalignments.
+Together, these factors likely contribute to the small but systematic clusters of elevated error observed in physiologically critical regions.
+These observations indicate that while the proposed framework captures global activation patterns and long-range structure effectively, resolving fine-scale, anatomy-driven variations remains challenging.
+By contrast, ventricular geometries parameterised using the UVC do not exhibit a comparable global seam-induced discontinuity.
+Instead, elevated errors are primarily confined to regions introduced by geometric truncation, such as the basal valve plane and the interface with the right ventricle.
+This contrast highlights that coordinate-induced limitations are strongly dependent on underlying anatomical topology rather than intrinsic to the operator-learning framework itself, and that intrinsic coordinate representations may behave differently across cardiac chambers.
 Finally, the framework was trained exclusively on high-fidelity simulated datasets.
-Although these simulations spanned a wide range of anatomies and electrophysiological parameters, direct validation against clinical electroanatomic mapping data is essential to confirm robustness under real-world conditions.
-The current implementation focuses on the left atrium; extending the approach to whole-heart or biventricular modelling will require adapting the anatomical projection and scaling the data generation process.
-Future directions include combining hybrid simulation–clinical datasets, refining anatomical embeddings, and exploring adaptive spatial resolutions to capture fine-scale electrophysiological detail without compromising inference speed.
+Although these simulations span a wide range of anatomies and electrophysiological parameters, validation against clinical electroanatomic mapping data will be essential to assess robustness under real-world conditions.
+In addition, the current implementation focuses primarily on atrial and ventricular geometries in isolation; extending the approach to whole-heart or fully biventricular modelling will require adapting the anatomical parameterisation and scaling the data generation process.
+Future directions include integrating hybrid simulation–clinical datasets, refining anatomical representations in regions of high complexity, and exploring adaptive spatial resolution strategies to further improve local accuracy without compromising inference efficiency.
 4 Methods
-We developed a unified, scalable computational framework that overcomes the computational bottlenecks of conventional EP simulations (Fig.
-3).
-This framework is designed to circumvent the computational bottleneck inherent in conventional EP simulations.
-It leverages a diverse, high-fidelity dataset generated by a custom high-performance FEM solver to enable the training of a neural operator that can predict LAT maps with accuracy comparable to ground truth, but within milliseconds.
+We describe a computational framework that combines high-fidelity electrophysiology simulations with neural operator learning to enable efficient prediction of LAT maps (Fig. 4).
 Figure 4: The overview of our proposed computational framework for operator learning for electrophysiology (EP) simulations.
@@
 They are artefacts of data preparation, yet they change the geometric domain on which the operator is defined.
-4.1.2 UAC projection
-We projected the spatial coordinates and fibre directions on the mesh, along with the resulting LAT map, to structured grids of fixed size by UAC.
-The UAC system [uac] defines a bijective transformation
-$\Phi:\mathcal{M}\rightarrow[0,1]^{2},$
-where $\mathcal{M}\subset\mathbb{R}^{3}$ denotes the atrial surface manifold, and $0,1]^{2}$ is a normalised two-dimensional coordinate space.
-The map $\Phi$ is constructed using anatomical landmarks, typically the pulmonary vein junctions and the mitral annulus, as anchor points to define physiologically meaningful coordinates.
-These coordinates correspond to a geodesic distance-based unfolding of the atrial surface, such that anatomical regions (e.g., anterior wall, posterior wall, roof) consistently map to aligned regions of the 2D domain across patients.
-The standardised representation allows to map topological entities from different anatomy to corresponding regions.
-Let $f:\mathcal{M}\rightarrow\mathbb{R}$ be a scalar function defined on the atrial mesh (e.g., LAT values), and let $(\alpha_{i},\beta_{i})=\Phi(\mathbf{x}_{i})$ be the UAC coordinates of a mesh vertex $\mathbf{x}_{i}$.
-We then construct a function $\tilde{f}:[0,1]^{2}\rightarrow\mathbb{R}$ via spatial interpolation:
-$\tilde{f}(\alpha,\beta)=\mathcal{I}\left(\left\{(\alpha_{i},\beta_{i},f_{i})\right\}_{i=1}^{N}\right),$
-where $\mathcal{I}(\cdot)$ denotes a suitable interpolation operator.
-In practice, we use a hybrid of linear and nearest-neighbour interpolation to preserve continuity in densely sampled regions while maintaining robustness in sparsely sampled areas.
-The 3D vertex coordinates $(x,y,z)$ for a mesh are projected directly to the 2D grid of shape $(3,50,50)$ where the first dimension (3) represents the channels corresponding to the $x$, $y$, and $z$ coordinates, and the latter two dimensions ($50\times 50$) define the resolution of the structured grid in the UAC system.
-Fibre orientations, however, are originally defined per mesh element rather than per vertex.
-To obtain vertex-wise fibre directions suitable for UAC projection, we convert element-level vectors to nodal vectors using area-weighted averaging.
-For each vertex, we accumulate the fibre vectors from all incident triangles, scaling each vector by the area of its corresponding triangle so that larger elements contribute proportionally.
-The resulting vectors are normalised to unit length to preserve directional information.
-These vertex-level fibre directions are then mapped to the 2D grid using the same UAC interpolation procedure as these coordinates, resulting in a 2D grid of shape $(3,50,50)$.
+4.1.2 Universal coordinates projection
+Spatial coordinates, fibre directions, and LAT fields from patient-specific mesh geometries are projected onto structured grids of fixed resolution using universal anatomical coordinate systems.
+For the left atrium, UAC defines a bijective transformation
+$\Phi_{\mathrm{LA}}:\mathcal{M}\rightarrow[0,1]^{2},$
+where $\mathcal{M}\subset\mathbb{R}^{3}$ denotes the atrial surface manifold.
+The mapping is constructed from anatomical landmarks, including the pulmonary vein junctions and mitral annulus, yielding a geodesic distance-based unfolding that aligns corresponding anatomical regions across patients [uac].
+An analogous intrinsic parameterisation is applied to ventricular geometries using the UVC system, which maps the three-dimensional ventricular volume to a structured coordinate domain.
+Let $f:\mathcal{M}\rightarrow\mathbb{R}$ denote a scalar field on the mesh (e.g., LAT), and let $\boldsymbol{\xi}_{i}=\Phi(\mathbf{x}_{i})$ be the corresponding UAC or UVC coordinates of mesh point $\mathbf{x}_{i}$.
+A coordinate-domain representation $\tilde{f}$ is constructed via spatial interpolation,
+$\tilde{f}(\boldsymbol{\xi})=\mathcal{I}\!\left(\{(\boldsymbol{\xi}_{i},f_{i})\}_{i=1}^{N}\right),$
+where $\mathcal{I}(\cdot)$ combines linear interpolation in densely sampled regions with nearest-neighbour interpolation near anatomical boundaries.
+For atrial meshes, three-dimensional vertex coordinates $(x,y,z)$ are mapped to a grid in the UAC domain.
+Fibre orientations, originally defined per element, are converted to vertex-wise vectors using area-weighted averaging and normalised to unit length before projection.
+An equivalent procedure is applied in the UVC domain for ventricular geometries.
+After neural operator inference in the coordinate domain, predictions can be mapped back to the original patient-specific geometry.
+For each mesh point $\mathbf{x}\in\mathcal{M}$ with coordinate $\boldsymbol{\xi}=\Phi(\mathbf{x})$, the predicted value is recovered by interpolating the learned field $\hat{f}(\boldsymbol{\xi})$.
 4.1.3 Identify pacing locations
@@
 The spatial distribution of these seven sites on a representative LA mesh is mapped onto the UAC coordinate system in Fig.
-4.
+5.
 Figure 5: Marking pacing sites on an LA.
@@
 4.2 GPU-based FEM EP simulation
-To generate a large-scale, comprehensive dataset for training, we simulated LA electrical propagation using the widely adopted monodomain equation [geselowitz1983bidomain] (further implementation details are provided in Supplementary Section 1).
+To generate a large-scale, comprehensive dataset for training, we simulated LA electrical propagation using the widely adopted monodomain equation [geselowitz1983bidomain] (further implementation details are provided in Supplementary Section 1, with an example of simulated wavefront propagation shown in Supplementary Fig.
+1 and benchmark verification results in Supplementary Fig.
+2).
 This biophysical model captures the electrical conduction through anisotropic cardiac tissue, balancing capacitive and ionic currents with diffusive current flow, governed by the longitudinal ($\sigma_{l}$) and transverse ($\sigma_{t}$) conductivities.
-The primary and clinically significant output extracted from these simulations is the Local Activation Time (LAT) map, which defines the time $t$ at which the transmembrane potential $V$ first crosses a predefined threshold ($V_{\text{th}}$, typically 0 mV) at any location $x$ in the cardiac domain:
+The primary and clinically significant output extracted from these simulations is the LAT map, which defines the time $t$ at which the transmembrane potential $V$ first crosses a predefined threshold ($V_{\text{th}}$, typically 0 mV) at any location $x$ in the cardiac domain:
 $\mathrm{LAT}(x)=\min\left\{t>0\ \middle|\ V(x,t)>V_{\text{th}}\right\},$
@@
 These values imply corresponding $\sigma_{t}$ in the range 0.01–0.1 S/m.
-Each simulation was run for 600 ms to capture the full activation sequence, and the local activation time (LAT) map was recorded during runtime.
+Each simulation was run for 600 ms to capture the full activation sequence, and the LAT map was recorded during runtime.
 Each simulation required approximately one minute on an NVIDIA A100 GPU, enabling the generation of 210,000 simulations for Cohort A and 98,700 simulations for Cohort B, yielding a combined total of 308,700 simulations.
@@
 The model is designed to process heterogeneous inputs, including sparse point clouds (e.g., pacing coordinates) and spatially distributed scalar fields (e.g., tissue conductivity, surface area), by transforming them into a unified, fixed-resolution tensor representation suitable for transformer-based spatial modelling (Fig.
-5).
+6).
 Figure 6: Model architecture.
@@
 To encode pacing (stimulation) locations provided as sparse, unordered two-dimensional point clouds of shape $(N,2)$ where $N$ is the number of vertices in the pacing region, we introduce an embedding module that maps the point set into a 2D grid representation (Fig.
-5b).
+6b).
 This module follows the architectural design of PointNet [qi2017pointnet], applying a multi-layer perceptron (MLP) independently to each input point, followed by a permutation-invariant mean pooling operation to obtain a global descriptor of the stimulation pattern.
@@
 The input to the neural operator is a multi-modal, multi-channel grid that encodes patient-specific mesh properties and simulation parameters into a fixed-size ten-channel representation (Fig.
-5a), which can be divided into two groups.
+6a), which can be divided into two groups.
 The first group comprises the seven geometric and anatomical channels, including the UAC-projected 3D vertex coordinates $(x,y,z)$ and the 3D fibre direction vectors (six channels in total, each forming a $(3,50,50)$ feature map).
@@
 This sequence is then processed by a stack of standard ViT blocks, each comprising multi-head self-attention and feed-forward sublayers, which enables the integration of global context and the capture of long-range spatial interactions (Fig.
-5c).
+6c).
 Following the ViT-based encoder, the latent tokens are reshaped into a compact 2D feature map.
@@
 All code used to generate the results in this study is available at: https://github.com/sagebei/unified_framework_operator_learning.
+Author contributions
+B.Z.
+developed the computational methodology, performed simulations, and drafted the manuscript.
+C.C.
+contributed to model development and data analysis.
+S.Q.
+assisted with numerical experiments and model development.
+M.B.
+contributed to software implementation and optimisation.
+A.W.C.L.
+prepared the data.
+C.R.
+contributed to manuscript editing and provided the simulation data.
+C.R., M.J.W.G.
+and L.H.G.A.H.
+provided experimental data.
+M.Q.
+assisted with analysing results.
+S.N.
+supervised the study, contributed to conceptualisation, and revised the manuscript.
+All authors reviewed the manuscript.
+Competing interests
+The authors declare no competing interests.

### v3

@@
-A unified framework for geometry-independent operator learning in cardiac electrophysiology simulations
+A unified framework for geometry-independent operator learning in cardiac electrophysiology
 Abstract
-Learning biophysically accurate solution operators for cardiac electrophysiology is fundamentally challenged by geometric variability across patient-specific heart anatomies.
-Most existing neural operator approaches are limited to structured or weakly deformed domains, restricting their applicability to realistic atrial and ventricular geometries.
-Here, we introduce a unified operator-learning framework that projects inputs and outputs onto a standardised anatomical coordinate system, decoupling electrophysiological dynamics from mesh topology.
-This formulation enables geometry-independent learning while preserving physiologically meaningful spatial organisation, and allows predictions to be interpolated back onto patient-specific geometries for anatomical interpretation.
-To support large-scale training within the framework, we develop a GPU-accelerated electrophysiology solver and generate over 300,000 high-fidelity simulations across diverse patient-specific left atrial geometries with varied pacing and conduction properties.
-Within this anatomical coordinate domain, we design a neural operator to predict full-field local activation time maps, achieving a mean absolute error of 5.1 ms and an inference time of 0.12 ms per sample, outperforming existing operator learning and convolutional baselines.
-We further validate the framework on ventricular geometries, demonstrating robust generalisation beyond the atrial setting.
-Together, this framework establishes a scalable foundation for fast, geometry-invariant cardiac electrophysiology modelling, with potential relevance for real-time and population-scale clinical workflows.
+Learning neural operators on heterogeneous and irregular geometries remains a fundamental challenge, as existing approaches typically rely on structured discretisations or explicit mappings to a shared reference domain.
+We propose a unified framework for geometry-independent operator learning that reformulates the learning problem in an intrinsic coordinate space defined on the underlying manifold.
+By expressing both inputs and outputs in this shared coordinate domain, the framework decouples operator learning from mesh discretisation and geometric variability, while preserving meaningful spatial organisation and enabling faithful reconstruction on the original geometry.
+We demonstrate the framework on cardiac electrophysiology, a particularly challenging setting due to extreme anatomical variability across heart geometries.
+Leveraging a GPU-accelerated simulation pipeline, we generate large-scale datasets of high-fidelity electrophysiology simulations across diverse patient-specific anatomies and train customised neural operators to predict full-field local activation time maps.
+The proposed approach outperforms established neural operators on both atrial and ventricular geometries.
+Beyond cardiac electrophysiology, we further show that the same representation enables operator learning in cardiac biomechanics, a distinct problem involving volumetric deformation, highlighting the generality of the proposed framework.
+Together, these results establish intrinsic coordinate representations as a principled and extensible pathway for neural operator learning on complex physical systems characterised by heterogeneous geometry.
 1 Introduction
-Computational models of cardiac electrophysiology provide a principled framework for studying electrical wave propagation, investigating arrhythmia mechanisms, and informing patient-specific therapeutic interventions [niederer2019computational, trayanova2024computational, prakosa2018personalized].
-A central application of these models is the analysis of local activation time (LAT) maps, which characterise the spatiotemporal evolution of electrical activation across the myocardium [williams2018local, jaffery2024review, karoui2021cardiac].
-Generating and analysing such LAT maps typically requires large ensembles of high-fidelity finite element method (FEM) simulations on anatomically detailed, patient-specific geometries, involving the numerical solution of coupled partial differential equations governing cardiac electrophysiology [niederer2011verification, opencarp].
-Atrial fibrillation (AF) is a primary clinical setting in which such patient-specific LAT analyses are routinely required [sahli2018interpreting].
-As the most common sustained cardiac arrhythmia, AF affects approximately 46.3 million individuals worldwide [kornej2020epidemiology] and is a leading cause of stroke [cheng2024global], heart failure [li2025global], and sudden cardiac death [lai2025causes].
-Computational modelling has therefore become an important tool in AF research and clinical studies, supporting mechanistic investigation of arrhythmia drivers, evaluation of ablation strategies, and development of personalised treatment planning [heijman2021computational, boyle2019computationally, roney2021time, krogh2012effects].
-However, the substantial computational cost of FEM-based workflows, often requiring hours to days per simulation, limits their applicability in settings that demand rapid or real-time decision-making [niederer2011simulating, boyle2019computationally].
-Recent advances in machine learning have opened new avenues for accelerating cardiac electrophysiology simulations.
-Neural operators, in particular, aim to learn mappings between infinite-dimensional function spaces [kovachki2023neural], enabling solution fields to be predicted for unseen inputs, discretisations, or geometries without retraining.
-Architectures such as DeepONet [don] and the Fourier Neural Operator (FNO) [fno] have demonstrated strong performance on parametric partial differential equation problems.
-Despite this promise, most existing neural operator frameworks are developed for structured or weakly deformed computational domains, which limits their applicability to the highly irregular, patient-specific geometries encountered in cardiac modelling [geofno, gino].
-In addition, training robust and generalisable operators requires large, high-quality simulation datasets, which remain expensive to generate using conventional electrophysiology solvers [ahmad2024diffeomorphic, trayanova2021machine].
-Several strategies have been proposed to address anatomical variability in operator learning.
-Diffeomorphic mapping approaches [zhao2025diffeomorphism], including Diffeomorphic Mapping Operator Learning (DIMON) [yin2024scalable], map individual geometries to a shared reference domain, enabling learning in a common representation while encoding geometry as auxiliary information.
-While effective for relatively simple geometries, such as the semi-spherical left ventricle, these approaches become challenging for anatomies with complex topology, such as the multi-holed left atrium, where constructing globally consistent mappings is non-trivial.
-Related methods, such as Geo-FNO [geofno], deform irregular geometries to structured domains to enable spectral learning, but the accuracy and stability of such mappings may deteriorate for highly intricate anatomical structures.
-Here, we propose a framework that reformulates neural operator learning in a coordinate domain intrinsic to cardiac anatomy rather than in the space of mesh discretisations.
-By expressing model inputs and outputs in a shared anatomical parameterisation, electrophysiological patterns can be represented consistently across patients despite substantial geometric and topological variability.
-Predictions obtained in this coordinate space can subsequently be mapped back onto the original patient-specific geometry, enabling anatomically faithful reconstruction and direct clinical interpretation.
-In this work, we adopt the Universal Atrial Coordinates (UAC) [uac] and Universal Ventricular Coordinates (UVC) [uvc] systems to define such intrinsic representations for atrial and ventricular geometries, respectively.
+Learning solution operators for partial differential equations on complex geometries remains a fundamental challenge across many areas of computational science.
+Neural operators offer a promising paradigm by learning mappings between functional spaces, enabling solution fields to be predicted for unseen parameters, discretisations, or geometries without retraining [kovachki2023neural].
+Architectures such as DeepONet [don] and the Fourier Neural Operator (FNO) [fno] have shown strong performance on parametric PDE problems, particularly on structured or weakly deformed domains.
+However, extending these approaches to heterogeneous, irregular geometries remains difficult, as most neural operator formulations implicitly rely on consistent spatial discretisations or globally aligned coordinate systems [geofno, gino].
+A central obstacle is that real-world physical systems are often defined on manifolds with substantial variability in shape, topology, and mesh representation.
+When learning is performed directly in the space of mesh discretisations, geometric differences become entangled with the underlying physical solution, hindering generalisation across geometries and limiting the applicability of neural operators to patient-specific or instance-specific domains.
+Existing approaches address this challenge by deforming irregular geometries to reference domains or by encoding geometry as auxiliary input [yin2024scalable, zhao2025diffeomorphism].
+While effective for relatively simple shapes, such strategies can become unstable or inaccurate for domains with complex topology, multiple boundaries, or substantial variability across instances.
+Here, we propose a general framework for geometry-independent neural operator learning that reformulates the learning problem in a coordinate domain intrinsic to the underlying manifold, rather than in the space of mesh discretisations.
+By expressing model inputs and outputs in a shared intrinsic parameterisation, spatial solution fields can be represented consistently across heterogeneous geometries despite substantial variability in shape, topology, and discretisation.
+Learning is performed entirely in this intrinsic coordinate domain, while predictions are subsequently mapped back to the original physical geometry for faithful reconstruction and interpretation.
+Importantly, the framework itself only assumes the existence of a smooth intrinsic coordinate chart on the manifold, making it broadly applicable beyond any specific physical system.
+Cardiac electrophysiology serves as a challenging and clinically relevant testbed for this framework, due to extreme anatomical variability, complex topology, and the need for high spatial fidelity.
+Computational models of cardiac electrophysiology are widely used to study electrical wave propagation, investigate arrhythmia mechanisms, and inform patient-specific therapeutic interventions [niederer2019computational, trayanova2024computational, prakosa2018personalized].
+A key output of such models is the local activation time (LAT) map, which characterises the spatiotemporal evolution of electrical activation across the myocardium [williams2018local, jaffery2024review, karoui2021cardiac].
+Generating LAT maps typically requires large ensembles of high-fidelity finite element simulations on anatomically detailed, patient-specific meshes, making conventional workflows computationally expensive and difficult to scale [niederer2011verification, opencarp].
+In this work, we adopt the Universal Atrial Coordinates (UAC) [uac] and Universal Ventricular Coordinates (UVC) [uvc] as concrete instantiations of intrinsic coordinate representations for atrial and ventricular geometries, respectively.
 These coordinate systems preserve physiologically meaningful spatial organisation while providing a unified domain across diverse anatomies, enabling neural operators to learn spatially coherent solution mappings without sensitivity to irregular mesh connectivity or topological inconsistencies.
-To support operator training within this framework, we further develop a GPU-accelerated electrophysiology simulation pipeline that enables efficient generation of high-fidelity LAT maps at scale across diverse patient-specific geometries.
-By integrating high-performance simulation with neural operator learning in anatomical coordinate space, our approach addresses both the computational cost of cardiac electrophysiology modelling and the geometric variability inherent in patient-specific anatomies, providing a scalable pathway toward rapid, physiology-informed inference in atrial fibrillation.
+To support operator learning at scale, we further introduce an efficient GPU-accelerated electrophysiology simulation pipeline that enables the generation of large ensembles of biophysically detailed LAT maps across diverse patient-specific geometries.
+By integrating intrinsic coordinate representations with high-performance simulation and neural operator learning, the proposed framework addresses both the geometric heterogeneity and the computational cost that have limited prior approaches.
+While cardiac electrophysiology serves as a challenging and clinically relevant demonstration, the framework is broadly applicable to other physical systems in which intrinsic coordinates are available, illustrating a general pathway toward geometry-independent operator learning on complex manifolds.
 2 Results
-2.1 Large-scale EP simulations on GPUs
+2.1 Large-scale electrophysiology datasets in intrinsic coordinate space
+We first assess whether an intrinsic coordinate representation yields a consistent learning domain across heterogeneous geometries, using LAT map prediction in cardiac electrophysiology as a representative operator learning task.
 We generated a comprehensive set of biophysically detailed EP simulations using a custom GPU-accelerated FEM solver applied to two independent cohorts of AF patients (Cohort A: $n=100$; Cohort B: $n=47$).
 This process constructed two large-scale datasets, Dataset A and Dataset B, respectively (see Methods).
-To ensure robustness across diverse anatomical and physiological conditions, for each of the $147$ patient-specific LA, we simulated LAT maps from seven distinct pacing sites across the LA.
-At each site, we sampled $300$ combinations of the key conduction parameters, the longitudinal conductivity ($\sigma_{l}$) and the anisotropy ratio ($\sigma_{l}/\sigma_{t}$), using Latin hypercube sampling [mckay2000comparison] to cover a wide range of clinically relevant tissue properties.
+To ensure robustness across diverse anatomical and physiological conditions, for each of the $147$ patient-specific left atriums (LAs), we simulated LAT maps from seven distinct pacing sites across the LA.
+At each site, we sampled $300$ combinations of the key conduction parameters, the longitudinal conductivity ($\sigma_{l}$) and the anisotropy ratio ($\sigma_{l}/\sigma_{t}$), using Latin Hypercube Sampling (LHS) [mckay2000comparison] to cover a wide range of clinically relevant tissue properties.
 This yields a total of 308,700 simulated LAT maps, establishing one of the largest computational datasets of its kind for training neural operators.
@@
 a, Example local activation time (LAT) map from a 3D left atrium (LA) electrophysiology simulation with coronary sinus pacing, together with its projection onto the standardised Universal Atrial Coordinate (UAC) domain.
-b, UMAP embedding of all UAC-projected LAT maps from Datasets A and B, showing seven pacing-site clusters per cohort and systematic displacement between corresponding clusters, indicating a pronounced domain shift between the two datasets.
+b, UMAP embedding of all UAC-projected LAT maps from Datasets A and B, showing seven clusters per cohort corresponding to pacing sites and systematic displacement between corresponding clusters, indicating a pronounced domain shift between the two datasets.
 c, Distribution of maximum LAT across pacing locations in Dataset A, demonstrating linear dependence on LA surface area and sensitivity to conductivity.
@@
 1a illustrates an LAT map on the 3D atrial surface together with its corresponding 2D UAC projection, with the coronary sinus pacing site producing full atrial activation within 204 ms.
-To assess potential cohort-level differences, we performed dimensionality reduction on the full set of UAC-projected LAT maps.
+To characterise variability within this intrinsic representation, we performed dimensionality reduction on the full set of UAC-projected LAT maps.
 High-dimensional LAT fields were first denoised using principal component analysis (PCA), and the resulting principal components were subsequently embedded using Uniform Manifold Approximation and Projection (UMAP) [healy2024uniform].
@@
 These consistent trends confirm that surface area and conductivity jointly modulate global activation timing, motivating their inclusion as key anatomical features in the subsequent analyses.
-2.2 Input feature sensitivity analysis
+2.2 Sensitivity to intrinsic representation and input features
 Figure 2: Comprehensive evaluation of model performance for LAT map prediction.
@@
 The notation $X\rightarrow Y$ indicates that the model was trained on source dataset $X$ and evaluated on target dataset $Y$.
-We quantified the contribution of individual anatomical input modalities by systematically removing spatial coordinates, fibre orientation, and surface area features, retraining the model under each configuration.
+We quantified the contribution of individual anatomical input modalities by removing spatial coordinates, fibre orientation, and surface area features, retraining the model under each configuration.
 Model performance was evaluated using mean absolute error (MAE), which captures numerical accuracy in milliseconds, and the structural similarity index measure (SSIM) [wang2004image], which assesses spatial fidelity of the predicted activation maps.
@@
 In contrast, removing fibre orientation features unexpectedly improves predictive performance across both metrics.
-This observation indicates that, in the present formulation, fibre orientation does not provide an additional informative signal beyond that already captured by geometric coordinates and global anatomical features.
-Based on this result, fibre orientation inputs were excluded from all subsequent experiments.
-Explanations for this effect are discussed further below.
-2.3 Evaluating the effect of regularisation
+This observation indicates that, in the present formulation, fibre orientation does not provide an informative signal beyond that already captured by geometric coordinates and global anatomical features.
+Consequently, fibre orientation inputs were excluded from all subsequent experiments.
+This finding does not imply that fibre orientation is unimportant for cardiac electrophysiology, but rather that its current atlas-based representation, when projected into the intrinsic coordinate domain, does not contribute additional predictive value for this task.
+2.3 Effect of spatial regularisation on operator learning
 We investigated the impact of spatial regularisation on LAT map reconstruction by augmenting the $H^{1}$ loss with two commonly used smoothness terms: total variation (TV) and a second-order Laplacian penalty, applied individually or in combination.
@@
 In contrast, applying either TV or Laplacian regularisation alone resulted in smaller and less consistent improvements, suggesting that their combined effect is necessary to adequately capture the spatial organisation of activation patterns.
-2.4 Benchmarking model performance
+2.4 Comparison with existing neural operator architectures
 We benchmarked our approach against a range of neural operators and convolutional architectures commonly used for modelling cardiac activation from multimodal inputs.
@@
 On a modern NVIDIA GPU, the model achieved an average inference time of approximately $0.12$ ms per sample.
-This represents orders of magnitude acceleration relative to conventional finite-element electrophysiology solvers, enabling near real-time LAT prediction across diverse anatomies and pacing configurations.
-2.5 Model generalisation across pacing sites
+This represents orders of magnitude acceleration relative to conventional FEM electrophysiology solvers, enabling near real-time LAT prediction across diverse anatomies and pacing configurations.
+2.5 Generalisation across stimulation conditions
 Generalisation across diverse pacing conditions is critical for scalable and clinically useful prediction of atrial activation, as it determines whether a single model can match or exceed the performance of models trained for individual pacing sites.
@@
 2e, right), reflecting superior preservation of spatial structure and clinically relevant activation patterns.
-2.6 The challenge of latent data heterogeneity
+2.6 Generalisation under latent domain shifts
 Our initial experiments confirmed the strong performance of our model on a single dataset (Dataset A), demonstrating its ability to capture anatomical features relevant to the target task within a controlled domain.
@@
 Training on the combined datasets $(A+B)$ encourages the model to focus on consistent anatomical features, ignoring spurious correlations tied to individual acquisition or annotation pipelines.
-Across both datasets, MAE decreased, and SSIM increased in comparison with the internal performance, demonstrating that exposure to independent cohorts acts as a form of regularisation.
+Across both datasets, MAE decreased, and SSIM increased in comparison with the internal performance, demonstrating that exposure to independent cohorts acts as a form of regularisation and that geometry-independent representations alone are insufficient for translational robustness unless coupled with heterogeneous, multi-domain training.
 For completeness, we also evaluated the model trained and tested on the combined datasets $(A+B)\rightarrow(A+B)$, obtaining $(5.77\pm 0.16)$ for MAE and $(0.9734\pm 0.003)$ for SSIM.
-2.7 Interpolation to patient-specific geometry
+2.7 Reconstruction on original geometries
 To assess anatomical prediction fidelity after interpolation, we evaluated model performance on both LA surface meshes and left ventricular (LV) volume meshes (Fig.
@@
 Model performance was evaluated using a previously proposed benchmark task in which LAT is predicted from one of seven pacing locations alone [yin2024scalable].
-Our experiments were conducted on Cohort A using 100 left atrial geometries, of which 20 were held out for testing.
-In addition, 1,006 left ventricular (LV) meshes were obtained from a publicly available dataset [lvdata], with 106 meshes reserved for evaluation.
+Our experiments were conducted on Cohort A using 100 LA geometries, of which 20 were held out for testing.
+In addition, 1,006 LV meshes were obtained from a publicly available dataset [lvdata], with 106 meshes reserved for evaluation.
 Model inputs and outputs were represented on a $100\times 100$ grid for LAs and a $50\times 50\times 50$ volumetric grid for LVs.
@@
 These ventricular error levels improve upon those reported in prior geometry-independent operator learning approaches, where mean absolute errors of approximately 5–6 ms were observed for LV meshes [yin2024scalable].
+The comparison is based on matched dataset scale, task definition, and activation time range, rather than a direct reimplementation, which is not feasible because the prior work relies on a simplified LV representation that is not compatible with the extracted LV meshes considered here.
 In both atrial and ventricular settings, errors remain consistent across pacing locations, indicating stable performance independent of stimulation site.
@@
 For the LV, larger errors are observed near the basal openings, corresponding to truncation at the valve plane, and along the interface with the right ventricle, reflecting geometric discontinuities introduced during extraction from bi-ventricular meshes.
+2.8 Application to kinematic deformation modelling
+To showcase that the proposed approach is not specific to cardiac electrophysiology, we apply it to a numerically distinct problem: kinematic deformation of non-linear elastic solids under load.
+This task involves volumetric deformation governed by material stiffness and applied pressure, and therefore represents a fundamentally different class of physical system and boundary conditions.
+We generate a dataset of three-dimensional geometries by varying two independent geometric parameters controlling global curvature and wall thickness.
+For each geometry, kinematic deformation simulations are performed across different combinations of material stiffness and internal pressure under fixed boundary constraints.
+As in the cardiac electrophysiology setting, a problem-specific intrinsic coordinate system is constructed to provide a consistent parameterisation across heterogeneous meshes, enabling all displacement fields to be represented in a shared coordinate domain.
+The neural operator takes four scalar inputs describing geometry and loading conditions and predicts the full-field displacement magnitude.
+Performance is evaluated using ten-fold cross-validation.
+The average displacement magnitude in the dataset is approximately $2\,\mathrm{mm}$, with a maximum of about $7\,\mathrm{mm}$, while the model achieves a mean absolute error of $0.35\pm 0.28\,\mathrm{mm}$ across test folds.
+Relative to the characteristic displacement magnitudes in the dataset, this error level indicates that the learned operator captures the dominant deformation behaviour across heterogeneous geometries.
 3 Discussion
 3.1 Geometry-independent neural operator learning
-Our approach demonstrates that a well-designed coordinate system is an efficient tool for bridging the gap between irregular, real-world data and the structured-domain requirements of neural networks, allowing a single neural operator to learn a geometry-independent mapping.
-By leveraging the universal coordinate system, we established a canonical representation that enabled neural operators to robustly capture the complex, nonlinear physics of cardiac conduction [hao2023gnot, li2023transformer, shih2025transformers, ovadia2024vito].
-The resulting prediction accuracy, with mean absolute errors on the order of 5 ms, lies within the typical measurement uncertainty and physiological variability reported for clinical electroanatomic mapping systems [gaeta2021high] (typically on the order of 5–10 ms [narayan2024advanced]).
-This demonstrates that our model’s predictions are not only numerically accurate but also fall within a clinically acceptable margin.
-A central premise of this work is that effective neural operator learning in cardiac electrophysiology depends on access to large, diverse, and high-quality training data.
-To this end, we developed a GPU-accelerated finite element electrophysiology solver that enables efficient generation of high-fidelity simulations across a wide range of patient-specific left atrial anatomies and electrophysiological conditions.
-This data generation pipeline addresses a longstanding computational bottleneck in cardiac modelling and provides a scalable foundation for learning generalisable electrophysiological structure.
-Beyond supporting our framework, this large, curated dataset provides a foundation for future data-driven studies in cardiac electrophysiology.
-Its scale, diversity, and consistency enable predictive models to generalise across anatomies and protocols, a capability previously limited by data scarcity.
-For context, G-FuNK [loeffler2024graph] was trained on 25 left atrial geometries, while DIMON [yin2024scalable] leveraged 1,006 ventricular geometries simulated at multiple pacing sites.
-In contrast, our simulations preserve these locations while also varying tissue conductivity, allowing neural operators to capture electrophysiological responses across diverse stimulation patterns and heterogeneous, patient-specific left atrial anatomies.
-Overall, our dataset spans 308,700 simulations, offering a far richer foundation for training and benchmarking state-of-the-art neural operators.
-The observation that removing fibre orientation features improved model performance was initially counterintuitive.
-Our analysis suggests two primary reasons for this finding.
-First, the current representation of fibre orientation may contain inherent noise or inaccuracies that interfere with the model’s ability to learn the true underlying electrophysiological dynamics.
-Second, the model appears to be leveraging the spatial and geometric priors provided by the coordinates and surface area features more effectively.
-These geometric inputs may implicitly contain sufficient information about the overall anatomical structure, making the explicit fibre orientation data redundant or even counterproductive in its current form.
-3.2 Comparative model analysis
-Our model achieves the best performance, which can be attributed to two key factors.
-First, the self-attention mechanism enables global context modelling [vaswani2017attention], allowing the network to capture long-range dependencies such as conduction wavefront spread from distant pacing sites.
-Second, by using convolutional patch embeddings and transposed convolution decoding, ViT maintains strong spatial localisation [wang2022convolutional], which is critical for resolving the high-frequency components of activation fronts.
-This balance between global awareness and local detail explains ViT’s superior pointwise accuracy and its strong preservation of spatial gradients.
-We benchmarked our approach against several state-of-the-art neural operators and two different CNN architectures.
-Our model’s performance surpassed these baselines for reasons rooted in their fundamental design.
-For instance, DeepONet, which in this work employs our model for its branch or trunk networks, leverages the same global context modelling advantages as our model.
-This enables DeepONet to perform reasonably well, particularly in SSIM and $H^{1}$ loss, by effectively capturing smooth functional mappings and spatial dependencies.
-However, despite incorporating ViT components, DeepONet’s architecture focuses on decomposing input-output relationships into branch and trunk networks, which may limit its end-to-end spatial locality learning compared to the full ViT model.
-Consequently, its MAE performance lags behind ViT, reflecting some difficulty in the precise localisation of rapid feature changes critical for accurate cardiac wavefront propagation.
-In contrast, FNOs are designed to learn global mappings in function space using spectral convolutions.
-While FNOs can capture global trends efficiently, their reliance on low-frequency Fourier modes makes them ill-suited for capturing sharp transitions and local gradient variations common in LAT maps, particularly near pacing initiation sites.
-This is reflected in their relatively poor $H^{1}$ loss and MAE, despite modest SSIM performance.
-WNOs attempt to address this by incorporating multi-resolution representations, theoretically improving sensitivity to localised patterns.
-However, WNO’s performance degrades substantially in this setting, likely due to instability in wavelet decomposition on structured but irregular biomedical domains and sensitivity to small-scale alignment errors, leading to high variance in $H^{1}$ loss.
-Among the convolutional baselines, U-Net delivers strong SSIM, indicating that it generates visually plausible LAT maps.
-Its hierarchical multi-scale architecture with skip connections allows for reconstruction of structural details.
-However, convolutional models like U-Net and ResNet-style CNN architectures are inherently local in receptive field, making it difficult to capture global pacing effects or subtle long-range interactions without very deep networks or large kernels [luo2016understanding].
+A key insight arising from this work is that geometry independence in operator learning can be most effectively addressed at the level of representation rather than architecture.
+Within this framework, the role of the intrinsic coordinate system is not to simplify the underlying physics, but to provide a canonical representation in which spatially coherent learning becomes feasible.
+By embedding geometric structure directly into the learning representation, neural operators can exploit both local spatial relationships and global contextual information without being confounded by irregular mesh connectivity or patient-specific discretisation artefacts.
+Importantly, the framework itself is agnostic to the specific physical system, requiring only the availability of a smooth intrinsic coordinate chart that is consistent across instances.
+Potential applications extend to include fluid–structure interaction, geophysical modelling on planetary surfaces, and biological transport processes on irregular anatomical domains.
+Cardiac electrophysiology provides a particularly stringent and clinically relevant testbed for this methodology, due to the pronounced variability in anatomy, topology, and boundary conditions across patients.
+By leveraging universal anatomical coordinate systems, we instantiate the proposed framework in this setting and demonstrate that neural operators can capture the complex, nonlinear dynamics of electrical conduction across highly heterogeneous cardiac geometries.
+The resulting prediction accuracy, with mean absolute errors on the order of 5 ms, lies within the typical measurement uncertainty and physiological variability reported for clinical electroanatomic mapping systems [gaeta2021high] (typically 5–10 ms [narayan2024advanced]).
+This indicates that the learned operators are not only numerically accurate but also operate within a clinically acceptable margin.
+A further premise underlying geometry-independent operator learning is the need for large, diverse, and high-quality training data.
+To support this, we developed a pipeline featuring a GPU-accelerated finite element electrophysiology solver that enables efficient generation of high-fidelity simulations across a wide range of patient-specific left atrial anatomies and electrophysiological conditions.
+This data generation capability addresses a longstanding computational bottleneck in cardiac modelling and provides a scalable foundation for learning generalisable structure within the proposed framework.
+Beyond enabling the present study, the resulting dataset establishes a resource for future data-driven investigations.
+Its scale, diversity, and internal consistency allow predictive models to generalise across anatomies and simulation protocols, a capability that has previously been constrained by data scarcity.
+For comparison, G-FuNK [loeffler2024graph] was trained on 25 left atrial geometries, while DIMON [yin2024scalable] leveraged 1,006 ventricular geometries simulated at multiple pacing sites, yielding 7,042 simulation instances.
+In contrast, our simulations preserve multiple pacing locations while also varying tissue conductivity, allowing neural operators to capture electrophysiological responses across diverse stimulation patterns and heterogeneous, patient-specific anatomies.
+Overall, the dataset comprises 308,700 simulations, offering a substantially richer foundation for training and benchmarking state-of-the-art neural operators.
+3.2 Inductive biases for learning on heterogeneous geometries
+The comparative evaluation reveals that strong performance in geometry-independent operator learning depends less on the choice of a specific neural architecture than on the inductive biases it encodes.
+In particular, accurate reconstruction of spatial fields on heterogeneous domains requires a balance between global contextual reasoning and precise local spatial representation.
+Models that emphasise only one of these aspects consistently underperform in this setting.
+Our approach benefits from this balance by combining global context modelling with explicit spatial structure, enabling the network to capture long-range dependencies while preserving fine-scale localisation.
+This is particularly important for activation time prediction, where global wavefront propagation and sharp local gradients coexist within the same spatial field.
+Architectures that capture global dependencies without maintaining sufficient spatial fidelity struggle to resolve rapid local transitions, while those that prioritise local structure alone fail to model long-range interaction driven by wavefront propagation.
+These trends are consistent across the evaluated baselines.
+Operator learning methods designed to model smooth global mappings perform well in terms of structural similarity but exhibit reduced pointwise accuracy in regions with steep spatial gradients.
+Conversely, convolutional architectures with strong local inductive biases generate visually coherent predictions yet remain limited in their ability to account for global pacing effects without substantial increases in depth or receptive field.
+Multi-resolution approaches offer partial mitigation but remain sensitive to alignment errors and domain irregularities in this context.
+Together, these results indicate that geometry-independent operator learning on irregular anatomical domains places specific demands on model design.
+Architectures must simultaneously integrate global information and preserve spatial locality in order to achieve both numerical accuracy and structural fidelity.
+This requirement is largely independent of the specific application and is expected to generalise to other problems involving partial differential equations on complex geometries.
 3.3 Toward domain-invariant generalisation
-Prior studies have evaluated model performance primarily under variations in pacing site [yin2024scalable].
-In contrast, a central contribution of our work is demonstrating that a single model can generalise across both diverse pacing locations and simulation parameters.
-This capability is essential for clinical applicability, since the pacing site is often variable or unknown in practice.
-Our results also show that the unified model, trained jointly across all seven pacing sites, consistently and significantly outperforms models trained for individual pacing locations.
-The unified model achieves a 16.5% reduction in mean MAE, and a 1.6% increase in average SSIM.
-This demonstrates that our model effectively extracts generalizable electrophysiological principles that are independent of the specific pacing location, making it a robust and scalable solution for real-world clinical applications where prior knowledge of the stimulation site is unavailable.
-Our cross-domain experiments reveal a critical insight.
-The apparent limitation of single-domain models is not a shortcoming of the ViT architecture, but a reflection of latent heterogeneity in clinical imaging data.
-By explicitly combining datasets from distinct acquisition and annotation protocols, we demonstrate that the model naturally discards protocol-specific shortcuts and focuses on the fundamental anatomical structure.
-This approach not only rescues performance on external cohorts but also slightly improves internal predictions.
-This finding establishes a generalizable strategy for achieving domain-invariant medical image analysis.
-The universal coordinate projection provides the necessary geometry-invariant foundation, while training on heterogeneous data acts as a regularisation method in feature space, compelling the model to learn only the biophysically relevant structures common to both domains and emphasising that true translational power derives from integrating diverse, multi-centre data rather than from isolated single-site optimisation.
+Generalisation across heterogeneous input conditions is a central challenge for data-driven modelling on complex anatomical domains.
+Prior studies have primarily evaluated performance under variations in pacing site [yin2024scalable].
+In contrast, our results show that a single model can generalise across both diverse pacing locations and simulation parameters.
+Models trained jointly across all pacing sites consistently outperform those trained for individual locations, indicating that the learned representations capture electrophysiological structure that is not tied to a specific stimulation configuration.
+Beyond pacing variability, our cross-domain experiments reveal a more fundamental limitation.
+Performance degradation when models are trained on a single dataset and evaluated on an independent cohort is not primarily attributable to the choice of network architecture, but rather reflects latent heterogeneity in imaging, annotation, and preprocessing protocols.
+By explicitly combining datasets acquired under distinct conditions, the model is encouraged to discard protocol-specific shortcuts and instead focus on anatomical features that are consistent across domains.
+This strategy not only improves performance on external cohorts but also yields modest gains on internal test sets.
+Together, these observations suggest a general strategy for achieving domain-invariant learning in medical imaging and simulation-based modelling.
+Universal coordinate representations provide a geometry-invariant foundation, while training on heterogeneous data acts as an implicit regularisation in feature space, promoting representations that reflect underlying biophysical structure rather than dataset-specific artefacts.
+More broadly, these results indicate that robust translational performance is more likely to arise from integrating diverse, multi centre data than from optimising models within a single acquisition domain.
 3.4 Limitations and future directions
-Beyond feature representation, limitations also arise from the use of intrinsic anatomical coordinate systems.
-For the left atrium, the UAC system maps a closed, topologically complex three-dimensional surface onto a flat two-dimensional domain.
-This flattening necessarily requires introducing a cut line on the surface, which appears as a seam in the coordinate representation.
+One limitation revealed by our ablation studies concerns the role of fibre orientation features.
+Removing fibre orientation may initially appear counterintuitive, yet it consistently improved predictive performance.
+Two factors are likely to contribute to this outcome.
+First, the current fibre orientation representation may introduce noise or uncertainty that interferes with learning the underlying electrophysiological dynamics.
+Second, the model appears to rely more effectively on the spatial and geometric priors provided by the anatomical coordinates and surface area features, which may already encode sufficient structural information.
+In this setting, explicit fibre orientation inputs may therefore be redundant or even detrimental in their present form.
+Beyond feature representation, limitations also arise from the anatomical coordinate parameterisation itself.
+For the left atrium, the UAC system maps a closed, topologically complex three-dimensional surface onto a flat two-dimensional domain, which necessarily introduces a cut line that appears as a seam in the coordinate representation.
 As a result, points that are physically adjacent on the original anatomy become separated across this seam, creating an artificial spatial discontinuity for the learning model.
-Consistent with this construction, our results reveal a narrow, coherent band of elevated prediction error aligned with the UAC seam, indicating that the neural operator struggles to interpolate smoothly across this boundary.
-In addition to seam-related effects, further localised errors are observed in regions of pronounced anatomical and electrophysiological complexity, particularly around the pulmonary vein junctions and the left atrial appendage.
-These regions exhibit substantial inter-patient variability that cannot be fully normalised by the UAC projection, leading to residual geometric heterogeneity in the coordinate domain.
-Moreover, activation wavefronts in these areas often display steep local gradients in activation time, increasing sensitivity to small spatial misalignments.
-Together, these factors likely contribute to the small but systematic clusters of elevated error observed in physiologically critical regions.
-These observations indicate that while the proposed framework captures global activation patterns and long-range structure effectively, resolving fine-scale, anatomy-driven variations remains challenging.
+Consistent with this construction, our results reveal a narrow band of elevated prediction error aligned with the UAC seam, indicating difficulty in interpolating smoothly across this boundary.
+Additional localised errors are observed in regions of pronounced anatomical and electrophysiological complexity, particularly around the pulmonary vein junctions and the left atrial appendage.
+These regions exhibit interpatient variability that cannot be fully normalised by the UAC projection, leading to residual geometric heterogeneity in the coordinate domain.
+Together, these factors contribute to small but systematic clusters of elevated error in physiologically critical regions.
+While the proposed framework captures global activation patterns and structure over long spatial scales, resolving fine-scale anatomical variation remains challenging.
 By contrast, ventricular geometries parameterised using the UVC do not exhibit a comparable global seam-induced discontinuity.
-Instead, elevated errors are primarily confined to regions introduced by geometric truncation, such as the basal valve plane and the interface with the right ventricle.
-This contrast highlights that coordinate-induced limitations are strongly dependent on underlying anatomical topology rather than intrinsic to the operator-learning framework itself, and that intrinsic coordinate representations may behave differently across cardiac chambers.
+Instead, ventricular error behaviour is governed primarily by boundary effects arising from geometric truncation rather than by coordinate-induced artefacts.
+Importantly, these truncations preserve the overall ventricular topology and continuity of the domain.
+This contrast highlights that limitations associated with intrinsic coordinate representations depend on underlying anatomical topology rather than being intrinsic to the operator learning framework itself, and that coordinate-based representations may therefore behave differently across cardiac chambers.
 Finally, the framework was trained exclusively on high-fidelity simulated datasets.
 Although these simulations span a wide range of anatomies and electrophysiological parameters, validation against clinical electroanatomic mapping data will be essential to assess robustness under real-world conditions.
-In addition, the current implementation focuses primarily on atrial and ventricular geometries in isolation; extending the approach to whole-heart or fully biventricular modelling will require adapting the anatomical parameterisation and scaling the data generation process.
-Future directions include integrating hybrid simulation–clinical datasets, refining anatomical representations in regions of high complexity, and exploring adaptive spatial resolution strategies to further improve local accuracy without compromising inference efficiency.
+In addition, the current implementation focuses primarily on atrial and ventricular geometries in isolation.
+Extending the approach to whole-heart or fully biventricular modelling will require adapting the anatomical parameterisation and scaling the data generation process.
+Future directions include integrating hybrid simulation–clinical datasets, refining anatomical representations in regions of high complexity, and exploring adaptive spatial resolution strategies to improve local accuracy while preserving computational efficiency and physical fidelity.
 4 Methods
-We describe a computational framework that combines high-fidelity electrophysiology simulations with neural operator learning to enable efficient prediction of LAT maps (Fig. 4).
-Figure 4: The overview of our proposed computational framework for operator learning for electrophysiology (EP) simulations.
+4.1 Geometry-independent operator learning framework
+The proposed framework separates operator learning from geometric representation by introducing an intrinsic coordinate domain that serves as the sole space in which learning is performed.
+Both input fields and target outputs are represented on this intrinsic domain at a fixed resolution, which allows standard neural operator architectures defined on regular grids to be used directly.
+The original physical mesh is not used during learning itself and is used only to project quantities that depend on geometry and physical parameters into the intrinsic domain.
+The framework consists of three stages.
+First, a problem-specific intrinsic coordinate map is used to project geometric coordinates, boundary information, and physical parameters from a mesh domain to a structured coordinate grid.
+Second, high-fidelity simulations are performed on the original mesh, and the resulting output fields are projected to the same coordinate grid to form paired training data, which are then used to train a neural operator in the coordinate domain.
+Third, predictions produced on the coordinate grid are interpolated back to the original mesh to enable anatomically faithful visualisation and downstream analysis.
+In this study, cardiac electrophysiology serves as the primary use case, with the target field being the local activation time map (Fig.
+4).
+An additional kinematic deformation problem is used to demonstrate that the proposed framework can be applied unchanged to mechanically governed systems with different physics and outputs.
+Figure 4: Overview of the geometry-independent operator learning framework, shown for cardiac electrophysiology.
 a, Preparation of left atrium (LA) mesh, mapping the 3D vertex coordinates defining the mesh to a 2D grid by Universal Atrial Coordinates (UAC), and identifying the pacing locations.
-b, The application of our GPU-based FEM solver to produce a large amount of simulated local activation time (LAT) maps, which are also projected to 2D grids by UAC.
-c, The training of neural operators using the structured simulation data.
-4.1 Image data processing
-4.1.1 Datasets description
+b, High-fidelity simulation on the original mesh and projection of output fields to the same coordinate grid to form training pairs.
+c, Neural operator learning in the coordinate domain and reconstruction of predictions on the original mesh for evaluation and downstream analysis.
+4.2 Instantiation for cardiac electrophysiology
+4.2.1 Geometries and cohorts
 The full dataset comprised 147 patient-specific left LA anatomies derived from two large-scale clinical cohorts.
@@
 These differences are not physiological changes.
-These differences are not physiological.
 They are artefacts of data preparation, yet they change the geometric domain on which the operator is defined.
-4.1.2 Universal coordinates projection
-Spatial coordinates, fibre directions, and LAT fields from patient-specific mesh geometries are projected onto structured grids of fixed resolution using universal anatomical coordinate systems.
+4.2.2 Intrinsic coordinate projection and reconstruction
+The framework assumes the availability of an intrinsic coordinate mapping that provides a consistent parameterisation across heterogeneous geometries.
+Using such a mapping, spatial coordinates, fibre directions, and scalar fields defined on patient-specific meshes are projected onto structured grids of fixed resolution.
 For the left atrium, UAC defines a bijective transformation
@@
 For each mesh point $\mathbf{x}\in\mathcal{M}$ with coordinate $\boldsymbol{\xi}=\Phi(\mathbf{x})$, the predicted value is recovered by interpolating the learned field $\hat{f}(\boldsymbol{\xi})$.
-4.1.3 Identify pacing locations
+4.2.3 High-fidelity electrophysiology simulations
+To generate a large-scale, comprehensive dataset for training, we simulated LA electrical propagation using the widely adopted monodomain equation [geselowitz1983bidomain] (more implementation details are provided in Supplementary Section 1, with an example of simulated wavefront propagation shown in Supplementary Fig.
+1 and benchmark verification results in Supplementary Fig.
+2).
+This biophysical model captures the electrical conduction through anisotropic cardiac tissue, balancing capacitive and ionic currents with diffusive current flow, governed by the longitudinal ($\sigma_{l}$) and transverse ($\sigma_{t}$) conductivities.
 To ensure the generalisability of neural operators and to capture a wide range of conduction dynamics, the EP simulations were initiated using a distributed set of seven clinically relevant pacing sites in the LA: the Left Atrial Appendage (LAA), the Left Superior and Inferior Pulmonary Veins (LSPV, LIPV), the Right Superior and Inferior Pulmonary Veins (RSPV, RIPV), the Coronary Sinus (CS), and the Roof which is defined as the superior wall of the left atrium connecting the LSPV and RSPV [romero2017atrial, moreira2008atrial, elmariah2013coronary].
+Figure 5: Marking pacing sites on an LA.
+a, Anterior and posterior views of an LA mesh where the seven pacing locations have been marked.
+b, Spatial distribution of these pacing sites in UAC coordinates.
+The blue regions indicate the anatomical openings (pulmonary veins and mitral valve).
 Pacing locations are defined across all patient anatomies using the UAC system.
@@
 The stimulation region is subsequently defined by all mesh vertices located within a fixed Euclidean radius of $2.0\,\text{mm}$ from this central point, ensuring comparable stimulation volumes across all seven sites.
-The spatial distribution of these seven sites on a representative LA mesh is mapped onto the UAC coordinate system in Fig.
-5.
-Figure 5: Marking pacing sites on an LA.
-a, Anterior and posterior views of an LA mesh where the seven pacing locations have been marked.
-b, Spatial distribution of these pacing sites in UAC coordinates.
-The blue regions indicate the anatomical openings (pulmonary veins and mitral valve).
-4.2 GPU-based FEM EP simulation
-To generate a large-scale, comprehensive dataset for training, we simulated LA electrical propagation using the widely adopted monodomain equation [geselowitz1983bidomain] (further implementation details are provided in Supplementary Section 1, with an example of simulated wavefront propagation shown in Supplementary Fig.
-1 and benchmark verification results in Supplementary Fig.
-2).
-This biophysical model captures the electrical conduction through anisotropic cardiac tissue, balancing capacitive and ionic currents with diffusive current flow, governed by the longitudinal ($\sigma_{l}$) and transverse ($\sigma_{t}$) conductivities.
+The spatial distribution of these seven sites on a representative LA mesh is mapped onto the UAC coordinate system (Fig.
+5).
 The primary and clinically significant output extracted from these simulations is the LAT map, which defines the time $t$ at which the transmembrane potential $V$ first crosses a predefined threshold ($V_{\text{th}}$, typically 0 mV) at any location $x$ in the cardiac domain:
@@
 Executing these simulations across eight A100 GPUs reduced the total wall-clock time from an estimated 214 days (single GPU) to approximately 27 days.
-4.3 Neural operator learning
-We formulate LAT map prediction as a neural operator learning problem, in which the goal is to approximate a mapping from patient-specific anatomical and electrophysiological parameters to the corresponding spatial distribution of activation times.
+4.2.4 Operator learning
+Within the proposed framework, we formulate LAT map prediction as a neural operator learning problem, in which the goal is to approximate a mapping from patient-specific anatomical and electrophysiological parameters to the corresponding spatial distribution of activation times.
 Formally, we aim to learn an operator
 $\mathcal{G}:(A,\sigma,p)\mapsto\mathrm{LAT}(\mathbf{x}),$
-where $A$ denotes the atrial geometry, $\sigma$ the conductivity tensor, $p$ the pacing site location, and $\mathrm{LAT}(\mathbf{x})$ the activation time at position $\mathbf{x}\in\Omega$ where $\Omega$ is the computational c.
+where $A$ denotes the atrial geometry, $\sigma$ the conductivity tensor, $p$ the pacing site location, and $\mathrm{LAT}(\mathbf{x})$ the activation time at position $\mathbf{x}\in\Omega$ where $\Omega$ is the computational domain.
 This framework requires the generation of large and diverse training datasets that pair input configurations $(A,\sigma,p)$ with ground-truth LAT maps obtained from high-fidelity EP simulations.
-4.3.1 Model architecture
 Our proposed architecture integrates a coordinate embedding module for point cloud projection with a Vision Transformer-based encoder and a convolutional decoder, enabling high-fidelity spatial field prediction from both sparse and structured physiological data.
@@
 Unlike PointNet, which is optimised for classification and segmentation tasks, our adaptation is explicitly designed to produce spatially structured embeddings that are compatible with convolutional and transformer-based feature extractors.
-Multi-modal feature integration.
-The input to the neural operator is a multi-modal, multi-channel grid that encodes patient-specific mesh properties and simulation parameters into a fixed-size ten-channel representation (Fig.
-6a), which can be divided into two groups.
-The first group comprises the seven geometric and anatomical channels, including the UAC-projected 3D vertex coordinates $(x,y,z)$ and the 3D fibre direction vectors (six channels in total, each forming a $(3,50,50)$ feature map).
-This is supplemented by the total surface area of the left atrium, a global anatomical feature that is broadcast across the spatial domain to form a $(1,50,50)$ constant grid.
-The second group comprises the two parametric features including the two conduction velocities ($\sigma_{l}$ and $\sigma_{t}$), which are also broadcast across the spatial domain to form a $(2,50,50)$ feature map, and the encoded pacing location in a $(1,50,50)$ feature map.
-The output of the neural network is trained against the target LAT map of shape $(1,50,50)$.
-ViT-Based Spatial Representation Learning.
-To model spatial dependencies across the input domain, we employ a Vision Transformer architecture that encodes the multi-channel input grid into a sequence of contextualised latent representations.
-A convolutional patch embedding layer [wu2021cvt] first divides the input grid into non-overlapping patches and projects each patch into a shared embedding space, producing a sequence of latent tokens.
-Learnable positional embeddings are added to retain spatial structure within the tokenised representation.
-This sequence is then processed by a stack of standard ViT blocks, each comprising multi-head self-attention and feed-forward sublayers, which enables the integration of global context and the capture of long-range spatial interactions (Fig.
+Multimodal feature integration.
+The neural operator operates on a grid representation with multiple channels, which jointly encode mesh geometry and simulation parameters (Fig.
+6a).
+These channels can be grouped into two categories.
+The first category encodes geometric and anatomical information.
+It consists of the UAC-projected three-dimensional vertex coordinates $(x,y,z)$ and the corresponding fibre direction vectors, yielding six spatial feature maps of size $(50,50)$.
+In addition, the total surface area of the left atrium is included as a global anatomical descriptor and broadcast across the spatial domain to form a constant $(1,50,50)$ grid.
+The second category encodes simulation-specific parameters.
+This includes the longitudinal and transverse conduction velocities $(\sigma_{l},\sigma_{t})$, which are broadcast across the grid to form a $(2,50,50)$ feature map, as well as the pacing location, represented as a spatial distribution on a $(1,50,50)$ grid.
+The operator is trained to predict the corresponding LAT map, represented on a $(1,50,50)$ grid.
+Spatial Representation Learning.
+To model spatial dependencies within the intrinsic coordinate domain, we employ a Vision Transformer (ViT) architecture [vit] to encode the multi-channel input grid into contextualised latent representations.
+Attention-based models have demonstrated strong capability for learning solution operators of partial differential equations by capturing long-range interactions and global context [li2023transformer, wu2024Transolver].
+This capability is essential in our setting, where activation dynamics combine global wave propagation with sharp local transitions across the same spatial domain.
+The input grid is first processed by a convolutional patch embedding layer [wu2021cvt], which partitions the domain into non-overlapping patches and projects each patch into a shared embedding space, yielding a sequence of latent tokens.
+Learnable positional embeddings are added to preserve spatial ordering within the tokenised representation.
+The token sequence is then processed by a stack of standard ViT blocks, each composed of multi-head self-attention and feed-forward sublayers, enabling the integration of global context and the modelling of long-range spatial interactions (Fig.
 6c).
-Following the ViT-based encoder, the latent tokens are reshaped into a compact 2D feature map.
-To recover full spatial resolution, we employ a lightweight decoder consisting of a transposed convolution layer for upsampling, followed by a convolutional projection head that produces the final output map.
-The transposed convolution operation, originally introduced in deep learning by [zeiler2011adaptive] for adaptive deconvolutional networks and later popularised for CNN visualisation by [zeiler2014visualizing], enables learnable upsampling by applying the transpose of the convolution operator.
-This decoding step ensures that the global features encoded by the transformer are effectively translated back into a high-resolution spatial field for pixel-wise prediction.
-4.3.2 Sobolev Loss with Spatial Smoothness Regularisations
+Following the encoder, the latent tokens are reshaped into a compact two-dimensional feature map.
+To recover full spatial resolution, we employ a lightweight decoder consisting of a transposed convolution layer for upsampling, followed by a convolutional projection head that maps latent features to the output field.
+The transposed convolution operation, originally introduced by [zeiler2011adaptive] for adaptive deconvolutional networks and later popularised for convolutional network visualisation by [zeiler2014visualizing], enables learnable upsampling through the transpose of the convolution operator.
+This decoding stage ensures that the global representations learned by the transformer are translated back into a high-resolution spatial field suitable for pixel-wise prediction.
+Sobolev Loss with Spatial Smoothness Regularisations
 To ensure our model produces physiologically consistent LAT maps, we employ an objective function grounded in the relative Sobolev $H^{1}$ norm.
@@
 This combination ensures that the model’s predictions are both highly accurate and physiologically coherent.
+4.3 Instantiation for large deformation mechanics
+We consider kinematic deformation of non-linear elastic solids under load as a representative volumetric mechanics problem.
+The computational domain is defined as a truncated ellipsoidal shell, a canonical approximation historically used to model left ventricular geometry [van1980application].
+Although simplified, this geometry exhibits non-uniform curvature and provides a well-defined basal surface for the application of boundary conditions.
+Inflation simulations were performed by fixing all degrees of freedom at the basal plane and applying a linearly increasing internal pressure to the inner surface.
+The resulting displacement field is governed by wall thickness, material stiffness, and applied pressure.
+The material behaviour was modelled using a hyperelastic transversely isotropic formulation based on the Guccione constitutive law [guccione1991passive].
+Fibre directions were assigned circumferentially, with incompressibility enforced via a bulk modulus of $\kappa=1000,\mathrm{kPa}$.
+Fibre, cross-fibre, and fibre-sheet stiffness parameters were fixed at $b_{f}=b_{t}=b_{fs}=18.48$.
+A dataset of 1,000 geometries was generated by varying four parameters: inner cavity diameter, wall thickness, bulk stiffness, and maximum internal pressure.
+Parameter ranges are reported in Table 1.
+All simulations were performed using the Cardiac Arrhythmia Research Package (CARP) [vigmond2003computational], with each simulation executed on 32 CPU cores.
+For operator learning, displacement magnitude fields were projected from the physical mesh onto a shared intrinsic coordinate system defined by universal ventricular coordinates [bayer2018universal].
+All fields were represented on a fixed $50\times 50\times 50$ grid in this coordinate domain and mapped back to the original geometry for evaluation.
+Table 1: Inflation simulation parameter ranges.
+Parameter Range
+Inner cavity diameter (mm) 38–58 [lang2015recommendations]
+Wall thickness (mm) 4–14 [kawel2012normal]
+Bulk stiffness (dimensionless) 1–5
+Maximum internal pressure (kPa) 1–10
 Declarations